### v1

# On importance sampling and independent Metropolis–Hastings with an unbounded weight function

Abstract
Importance sampling and independent Metropolis–Hastings (IMH) are among the fundamental building blocks of Monte Carlo methods. Both require a proposal distribution that globally approximates the target distribution. The Radon–Nikodym derivative of the target distribution relative to the proposal is called the weight function. Under the weak assumption that the weight is unbounded but has a number of finite moments under the proposal distribution, we obtain new results on the approximation error of importance sampling and of the particle independent Metropolis–Hastings algorithm (PIMH), which includes IMH as a special case. For IMH and PIMH, we show that the common random numbers coupling is maximal. Using that coupling we derive bounds on the total variation distance of a PIMH chain to the target distribution. The bounds are sharp with respect to the number of particles and the number of iterations. Our results allow a formal comparison of the finite-time biases of importance sampling and IMH. We further consider bias removal techniques using couplings of PIMH, and provide conditions under which the resulting unbiased estimators have finite moments. We compare the asymptotic efficiency of regular and unbiased importance sampling estimators as the number of particles goes to infinity.
1 Introduction
1.1 Context and contributions
Monte Carlo with global proposals.
Monte Carlo methods aim to approximate a target distribution $\pi$ on a measurable space $(\mathbb{X},\mathcal{X})$, for example $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$. These techniques are crucial when analytical computation of expectations under $\pi$ is infeasible due to high dimensionality or complex distribution forms. The goal is to evaluate integrals of functions $f:\mathbb{X}\longrightarrow\mathbb{R}$ with respect to $\pi$:
[…]
Two primary approaches are Markov Chain Monte Carlo (MCMC) methods, that construct a Markov chain with $\pi$ as its stationary distribution, and Importance Sampling (IS) methods, where the target distribution is approximated by weighted samples. Among MCMC methods, the independent Metropolis–Hastings (IMH) algorithm is a specialized form of the Metropolis–Rosenbluth–Teller–Hastings (MRTH) algorithm (Metropolis et al., 1953; Hastings, 1970), in which proposals are drawn from a distribution $q$ independently of the current state of the chain. The same proposal $q$ can be employed to generate draws in importance sampling, a procedure that can be traced back to Kahn (1949), as noted in Andral (2022). We assume that we can draw from $q$ and that its density, also denoted by $q$, can be evaluated pointwise. Therefore, IMH and IS propose two ways of correcting for the discrepancy between a proposal $q$ and a target $\pi$, and their comparison is a natural and fundamental question.
Our contributions
concern the performance of importance sampling and independent Metropolis–Hastings, and their comparison. Our key assumption is that the Radon–Nykodym derivative $\omega$ of $\pi$ with respect to $q$, termed the weight, has $p$ finite moments under $q$. We first show that the bias of self-normalized importance sampling is of order $N^{-1}$, and we obtain new bounds on the moments of the error in importance sampling in Section 2. We then consider IMH, and show that the common random numbers coupling is optimal in Section 3. Using this coupling, in Section 4 we show that the total variation distance between IMH at iteration $t$ and $\pi$ decays as $t^{p-1}$. We obtain matching lower bounds in an example. We also obtain explicit dependencies in $N$ for the particle IMH algorithm (Andrieu et al., 2010). In Section 5 we consider the bias removal technique of Glynn & Rhee (2014) applied by Middleton et al. (2019) to the particle IMH algorithm. This yields an unbiased estimator that can be implemented whenever self-normalized importance sampling or IMH can be implemented. We provide conditions under which the estimators have finite moments, and conditions under which their efficiency is asymptotically equivalent to that of importance sampling.
1.2 Importance sampling
Self-normalized importance sampling (IS) is described in Algorithm 1, see also Chapter 9.2 in Owen (2013). Central to importance sampling is the weight function defined as
$\omega:x\mapsto\frac{\pi(x)}{q(x)},\quad\text{so that}\quad q(\omega)=1.$ (2)
Since multiplicative constants in $\omega$ have no effect on the IS estimator (4), it can be computed as long as the user can evaluate a function proportional to $\omega$ in (2). Unless specified otherwise, by IS we will refer to the self-normalized procedure in Algorithm 1; and not to the more basic estimator $N^{-1}\sum_{n=1}^{N}\omega(x_{n})f(x_{n})$ that depends on the multiplicative constant in $\omega$.
  1. 1.
    Sample $N$ particles independently $x_{1},\ldots,x_{N}$ from $q$.
  2. 2.
    Compute the importance weights $\omega(x_{n})=\pi(x_{n})/q(x_{n})$ for $n\in[N]=\{1,\cdots,N\}$.
  3. 3.
    Compute
    $\hat{Z}(x_{1},\ldots,x_{N})=N^{-1}\sum_{n=1}^{N}\omega(x_{n}).$ (3)
  4. 4.
    For any test function $f$, compute the IS estimator
    $\hat{F}(x_{1},\ldots,x_{N})=\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}.$ (4)
  5. 5.
    Return $\hat{F}(x_{1},\ldots,x_{N})$ and $\hat{Z}(x_{1},\ldots,x_{N})$.
Algorithm 1 Self-normalized importance sampling.
We make the following assumption throughout.
Assumption 1.
For any measurable set $A\in\mathcal{X}$, if $q(A)=0$, then $\pi(A)=0$, in other words $\pi$ is absolutely continuous with respect to $q$. Furthermore, $\omega(x)$ with $x\sim q$ is almost surely positive, and $q(\omega)=1$.
Under Assumption 1, if $\pi(f)$ exists then $\hat{F}(x_{1},\ldots,x_{N})\to\pi(f)$ as $N\to\infty$ almost surely. The asymptotic variance of IS is directly computed from the delta method (Owen, 2013; Robert & Casella, 2004; Liu, 2008), assuming $q(\omega^{2}\cdot f^{2})<\infty$ and $q(\omega^{2})<\infty$,
$\lim_{n\to\infty}\mathbb{V}\left[\sqrt{N}(\hat{F}(x_{1},\ldots,x_{N})-\pi(f))\right]=q(\omega^{2}\cdot(f-\pi(f))^{2}).$ (5)
Agapiou et al. (2017) provide non-asymptotic bounds on the mean squared error and on the bias of importance sampling, which are both inversely proportional to the number $N$ of draws from $q$; see Theorem 2.2. The exact form of the asymptotic bias of IS is well-known (e.g. Skare et al., 2003; Liu, 2008), and we provide a formal statement in Section 2.
1.3 Independent Metropolis–Hastings
Independent Metropolis-Hastings (IMH) is an instance of the Metropolis–Rosenbluth–Teller–Hastings algorithm, where the draws are proposed from $q$ independently of the current state of the chain (Hastings, 1970, Section 2.5); see Algorithm 2. Thus IMH is implementable in the same settings as importance sampling. Under Assumption 1, the IMH chain is $\pi$-irreducible, on top of being aperiodic and $\pi$-invariant by design, thus for $\pi$-almost every $x$, $|P^{t}(x,\cdot)-\pi|_{\mathrm{TV}}\to 0$ as $t\to\infty$ (Theorem 4 in Roberts & Rosenthal, 2004), where $P$ denotes the transition kernel of IMH, $P^{t}$ denotes the $t$-steps transition kernel, and $|\mu-\nu|_{\text{TV}}=\sup_{A\in\mathcal{X}}\mu(A)-\nu(A)$.
The asymptotic variance of the ergodic average $t^{-1}\sum_{s=0}^{t-1}f(x_{t})$ generated by IMH is finite if $\pi(f^{2})<\infty$ and $q(\omega^{2}\cdot f^{2})$ (Theorem 2 in Deligiannidis & Lee, 2018), and under these conditions its variance is greater than or equal to the asymptotic variance of importance sampling in (5) (Proposition 2 in Deligiannidis & Lee, 2018). Thus, in terms of asymptotic variance, the comparison is clear: IS outperforms IMH. Since IMH defines a Markov transition, it can directly be used as a step within an encompassing Gibbs sampler (Skare et al., 2003), and it is commonly used within sequential Monte Carlo samplers (Chopin, 2002; South et al., 2019), and thus has its specific uses irrespective of the performance comparison with importance sampling.
  1. 1.
    Draw $x^{\star}\sim q$.
  2. 2.
    Compute the acceptance probability:
    $\alpha_{\text{RH}}(x,x^{\star})=\min\left\{1,\frac{\omega(x^{\star})}{\omega(x)}\right\}.$ (6)
  3. 3.
    Draw $u$ from a $\text{Uniform}(0,1)$ distribution.
  4. 4.
    If $u<\alpha_{\text{RH}}(x,x^{\star})$, set $x^{\prime}=x^{\star}$, otherwise $x^{\prime}=x$.
  5. 5.
    Return $x^{\prime}$.
Algorithm 2 IMH algorithm describing one step starting from $x$.
When it comes to non-asymptotic behavior, for IMH there is an important distinction between two cases (Mengersen & Tweedie, 1996): either the weight is bounded, in which case the chain is geometrically ergodic and exact rates are obtained in Wang (2022), or the weight is unbounded and the convergence is polynomial at best; in the latter case, various results are provided e.g. in Jarner & Roberts (2002); Douc et al. (2007); Roberts & Rosenthal (2011); Andrieu et al. (2022) and Douc et al. (2018, Chapter 17); see Section 4.3. In Section 4 we provide polynomial bounds on the total variation distance to stationarity for IMH under moment conditions on $\omega$ under $q$. Our results enable a comparison of the biases of IS and IMH in Section 4.4, which turns out in favor of IMH.
In the following we consider the particle IMH (PIMH) generalization of IMH, where $N$ proposals are drawn at each iteration (Andrieu et al., 2010, Section 4.2); see Algorithm 3. We define the algorithm on the state space $\mathbb{X}^{N}$, use boldface to denote its elements, e.g. $\mathbf{x}=(x_{1},\ldots,x_{N})\in\mathbb{X}^{N}$, and denote the transition kernel by $P$. If $N=1$ the algorithm corresponds to IMH, and our results apply for all $N\geq 1$. To view Algorithm 3 as a special case of IMH, define for any $N\geq 1$
$\bar{\pi}(x_{1},\ldots,x_{N})$ $=\sum_{k=1}^{N}\frac{\pi(x_{k})}{N}\prod_{n\neq k}q(x_{n})=\left(\frac{1}{N}\sum_{k=1}^{N}\omega(x_{k})\right)\prod_{n=1}^{N}q(x_{n}),$ (7)
$\bar{q}(x_{1},\ldots,x_{N})$ $=\prod_{n=1}^{N}q(x_{n}),$ (8)
and, in the case $N=1$, $\bar{\pi}(x_{1})=\pi(x_{1})$. From (3) and the above definitions, we can write:
$\bar{\omega}(\mathbf{x})=\frac{\bar{\pi}(\mathbf{x})}{\bar{q}(\mathbf{x})}=\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n})=\hat{Z}(\mathbf{x}).$ (9)
Hence, IMH as in Algorithm 2, with proposal $\bar{q}$ and target $\bar{\pi}$, is equivalent to Algorithm 3.
  1. 1.
    Draw $\mathbf{x}^{\star}=(x^{\star}_{1},\ldots,x_{N}^{\star})\sim\bar{q}$.
  2. 2.
    Compute the acceptance probability:
    $\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})=\min\left\{1,\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})}\right\},$ (10)
    where $\hat{Z}:\mathbf{x}\mapsto N^{-1}\sum_{n=1}^{N}\omega(x_{n})$.
  3. 3.
    Draw $u$ from a $\text{Uniform}(0,1)$ distribution.
  4. 4.
    If $u<\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})$, set $\mathbf{x}^{\prime}=\mathbf{x}^{\star}$, otherwise $\mathbf{x}^{\prime}=\mathbf{x}$.
  5. 5.
    Return $\mathbf{x}^{\prime}$.
Algorithm 3 PIMH algorithm describing one step starting from $\mathbf{x}=(x_{1},\ldots,x_{N})$.
In order to estimate an expectation $\pi(f)$ from the PIMH output, a Rao–Blackwellization argument explained in Andrieu et al. (2010, Appendix B.5) leads to
$\mathbb{E}_{\mathbf{x}\sim\bar{\pi}}\left[\hat{F}(\mathbf{x})\right]=\mathbb{E}_{\bar{\pi}}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}\right]=\pi(f),$ (11)
with $\hat{F}$ as in (4). For completeness we restate the arguments of Andrieu et al. (2010) justifying (11) in Appendix A.1. Thus, we can compute $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{n=1}^{N}\omega(x_{n})$ at each iteration of the chain $(\mathbf{x}_{t})_{t\geq 0}$, and the ergodic average $T^{-1}\sum_{t=0}^{T-1}\hat{F}(\mathbf{x}_{t})$ may converge to $\pi(f)$. This allows connections between PIMH and IS, the latter being equal to $\hat{F}(\mathbf{x})$ with $\mathbf{x}\sim\bar{q}$.
1.4 Moment conditions on the weight
We introduce the assumption under which most of our results are derived.
Assumption 2.
The weights have a finite $p$-th moment for $p\geq 2$: $q(\omega^{p})<\infty$.
This is a weak and natural assumption in the context of both self-normalized importance sampling and IMH. For bounded test functions Assumption 2 is necessary for the asymptotic variance of both self-normalized importance sampling and IMH to be finite.
Example 1 (Exponential distributions).
Let $\pi$ be the Exponential(1) distribution and let $q$ be the Exponential($k$) distribution with $q(x)=ke^{-kx}$, both on $\mathbb{R}_{+}$. If $k\leq 1$, the weight $\omega(x)$ is upper bounded by $k^{-1}$, and Assumption 2 holds for all $p\geq 2$. If $k>1$, then $q(\omega^{p})<\infty$ holds with any $p<k/(k-1)$, and the requirement $p\geq 2$ translates into $k<2$. The example is considered in Jarner & Roberts (2007); Roberts & Rosenthal (2011); Andrieu et al. (2022).
Example 2 (Normal distributions).
Let $\pi$ be the Normal(0,1) distribution and let $q$ be the Normal$(0,\sigma^{2})$ distribution, both on $\mathbb{R}$. If $\sigma^{2}\geq 1$, the weight $\omega(x)$ is upper bounded by $\sigma$, and Assumption 2 holds for all $p\geq 2$. If $\sigma^{2}<1$, then $q(\omega^{p})<\infty$ holds for $p<\sigma^{-2}/(\sigma^{-2}-1)$. The requirement $p\geq 2$ in Assumption 2 translates into $\sigma^{2}>1/2$. The example is considered in Roberts & Rosenthal (2011); Owen (2013).
Assumption 2 implies the following well-known behavior of the average of $N$ independent weights. The proofs of the results below are in Appendix A.1.
Proposition 1.1.
Let $\mathbf{x}=(x_{n})_{n=1}^{N}$ be $N$ i.i.d. random variables from $q$. Under Assumptions 1-2, with $p\geq 2$, $\hat{Z}(\mathbf{x})=N^{-1}\sum_{n=1}^{N}\omega(x_{n})$ satisfies, for all $N\geq 1$,
$\mathbb{E}_{\bar{q}}[\hat{Z}(\mathbf{x})^{p}]$ $\leq\left(1+\frac{2^{1-1/p}(p-1)(1+q(\omega^{p}))^{1/p}}{\sqrt{N}}\right)^{p},$ (12)
$\mathbb{E}_{\bar{q}}[|\hat{Z}(\mathbf{x})-1|^{p}]$ $\leq\left(\frac{2^{1-1/p}(p-1)(1+q(\omega^{p}))^{1/p}}{\sqrt{N}}\right)^{p}=:M(p)N^{-p/2}.$ (13)
We will repeatedly use the following consequence of Markov’s inequality and Proposition 1.1.
Lemma 1.1.
Under Assumptions 1-2, with $p\geq 2$, there exists a constant $M(p)>0$ such that, for any $z>0$ and $N\geq 1$:
$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+z\right)\leq\frac{M(p)}{N^{p/2}z^{p}}.$ (14)
Remark 1.1.
Most of our proofs require that $\hat{Z}(\mathbf{x})$ is a non-negative random variable, with $\mathbb{E}_{\bar{q}}[\hat{Z}(\mathbf{x})]=1$, $\mathbb{E}_{\bar{q}}[|\hat{Z}(\mathbf{x})-1|^{p}]\leq M(p)N^{-p/2}$ for some $M(p)$ independent of $N$, but not directly that $\hat{Z}(\mathbf{x})$ is an average of i.i.d. weights. Thus $\hat{Z}(\mathbf{x})$ could for example be the normalizing constant estimator generated by a sequential Monte Carlo sampler. Results on moments of sequential Monte Carlo normalizing constant estimators can be found in e.g. Del Moral (2013, Section 16.5).
2 Bias and moments of importance sampling
Our first contribution is a clean statement on the asymptotic bias of self-normalized importance sampling. Introductory material on importance sampling often makes the point that the basic importance sampling estimator $N^{-1}\sum_{n=1}^{N}f(x_{n})\omega(x_{n})$ is unbiased, but since $\omega$ can only be evaluated up to a multiplicative constant, users may need to resort to the self-normalized estimator in (4), which is biased: $\mathbb{E}[\hat{F}(\mathbf{x})]\neq\pi(f)$. The form of the asymptotic bias is well known, e.g. Section 2.5. in Liu (2008). However, somewhat surprisingly, formal results appear to be lacking. The closest may be Theorem 2 in Skare et al. (2003), but their emphasis is on the pointwise relative error of the density of a particle selected from the IS approximation. Their Remark 1 translates this into a bound on the bias for bounded functions under the assumption of bounded weights. We provide Theorem 2.1, with a proof in Appendix A.2, which gives the leading term in the bias of IS under more general conditions on the weights.
Theorem 2.1.
Let $f$ with $\pi(|f|)<\infty$. Assume that $x_{1},\ldots,x_{n}$ are i.i.d. from $q$, let $\omega:x\mapsto\pi(x)/q(x)$, and let $\hat{F}(\mathbf{x})=\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/\sum_{m=1}^{N}\omega(x_{m})$. Assume that $q(|f|\cdot\omega^{2})<\infty$, $q(|f|\cdot\omega^{3})<\infty$ and $q(\omega^{-\eta})<\infty$ for some $\eta>0$. Then
$\lim_{N\to\infty}N\times\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[\hat{F}(\mathbf{x})-\pi(f)\right]$ $=-\int\left(f(x)-\pi(f)\right)\omega^{2}(x)q(dx).$ (15)
Theorem 2.1 assumes a finite inverse moment of the weight, and for bounded $f$ the theorem requires $q(\omega^{3})<\infty$. The inverse moment assumption may be removed at the cost of higher positive moments. Agapiou et al. (2017) provide an upper bound on the bias under weaker assumptions, which we restate below.
Theorem 2.2 (Bias part of Theorem 2.1 in Agapiou et al. (2017)).
Suppose that $q(\omega^{2})<\infty$ and that $|f|_{\infty}\leq 1$. Then, for all $N\geq 1$,
$\mathbb{E}_{\mathbf{x}\sim\bar{q}}[\hat{F}(\mathbf{x})-\pi(f)]\leq\frac{12}{N}q(\omega^{2}).$
Theorem 2.3 in Agapiou et al. (2017) provides upper bounds of order $N^{-1}$ also for unbounded test functions, under moment conditions on $f$ and on $f\cdot\omega$. Our Theorem 2.1 establishes that $N^{-1}$ is the exact order of the asymptotic bias as a function of $N$, but requires additional conditions. We next provide a result on the $s$-th moments of the error in importance sampling for unbounded test functions. Theorem 2.3 generalizes the MSE part of Theorem 2.3 in Agapiou et al. (2017) to arbitrary orders $s\geq 2$, and its assumptions are weaker in the case $s=2$, as discussed below. The proof is in Appendix A.2. The bounds are central to the results of Section 5.
Theorem 2.3.
Assume that there exist $p\in[2,\infty)$ and $r\in[2,\infty]$ such that $q(\omega^{p})<\infty$ and $q(|f|^{r})<\infty$, and $q(f^{2}\cdot\omega^{2})<\infty$, then for any $2\leq s\leq pr/(p+r+2)$ and any $N\geq 1$, we have:
$\mathbb{E}_{\bar{q}}\left[\left\lvert\hat{F}(\mathbf{x})-\pi(f)\right\rvert^{s}\right]\leq CN^{-s/2},$
where the constant $C$ depends on $r,p,s,q(|f|^{r}),q(\omega^{p}),q(f^{2}\cdot\omega^{2})$. When $r=\infty$, the statement holds for $f$ such that $|f|_{\infty}<\infty$ and all $s\leq p$.
A few remarks are in order:
  • The condition $s\leq pr/(p+r+2)$ implies $s\leq\min\{p,r\}$.
  • We have $q((f\omega)^{pr/p+r})<\infty$ if $q(\omega^{p})<\infty$ and $q(f^{r})<\infty$. Indeed, when $r<\infty$, $q((f\omega)^{pr/p+r})\leq q(f^{r})^{p/p+r}q(\omega^{p})^{r/p+r}<\infty$. When $r=\infty$, the claim remains correct (by understanding $pr/(p+r)$ as $p$), since $q((f\omega)^{p})\leq\lVert f\rVert_{\infty}^{p}q(\omega^{p})$. This observation leads to two facts: 1) If $pr/(p+r)\geq 2$ (e.g. $p=r=4$ or $p=2,r=\infty$), the assumption $q(f^{2}\cdot\omega^{2})<\infty$ in Theorem 2.3 can be derived from the assumptions $q(\omega^{p})<\infty$ and $q(f^{r})<\infty$. 2) The basic importance sampling estimator $N^{-1}\sum_{n=1}^{N}f(x_{n})\omega(x_{n})$ has a finite $s$-th moment under the same conditions, as it has a finite $pr/(p+r)$-th moment, and $s\leq pr/(p+r+2)\leq pr/(p+r)$.
  • We may be particularly interested in the mean-squared error (MSE) of IS, corresponding to $s=2$. Theorem 2.3 implies that the MSE is of order $1/N$ as long as $2\leq pr/(p+r+2)$. This condition holds, for example, if $\min\{p,r\}\geq 2(1+\sqrt{2})\approx 4.828$, or if $p\geq 3$ and $r\geq 10$, or if $p=2$ and $r=\infty$. The case $s=2$ can be compared to the MSE part of Theorem 2.3 in Agapiou et al. (2017). In our notation, they require $q(|f\cdot\omega|^{2d})<\infty$, $q(\omega^{2e})<\infty$, $q(|f|^{2a})<\infty$, $q(\omega^{2b(1+a^{-1})})<\infty$, for $a,b,d,e>1$ such that $a^{-1}+b^{-1}=1,d^{-1}+e^{-1}=1$. Their assumption implies ours, as can be seen by setting $r=2a$ and $p=2b(1+a^{-1})$, since then
    $\frac{pr}{(p+r+2)}=\frac{4b(a+1)}{2a+2b+2ba^{-1}+2}=\frac{2(a+1)}{a(1-a^{-1})+1+a^{-1}+(1-a^{-1})}=\frac{2(a+1)}{a+1}=2,$
    i.e. our theorem holds with $s=2$ under their assumptions.
3 Optimality of coupling IMH with common draws
With a view toward deriving upper bounds on the total variation distance of IMH to stationarity, we consider the common draws (or common random numbers) coupling of a generic IMH algorithm, described in Algorithm 4, and PIMH is retrieved as a special case. The coupling is very simple and was considered in Liu (1996); Roberts & Rosenthal (2011). The pseudocode describes the transition kernel $\bar{P}((\mathbf{x},\mathbf{y}),\cdot)$ of the coupled chains, and we denote the transition of IMH by $P$. It was remarked around Lemma 1 in Wang et al. (2021) that this coupling is “one-step maximal”, in the sense that the probability $\bar{P}((\mathbf{x},\mathbf{y}),D)$ where $D=\{(\mathbf{x},\mathbf{y}):\mathbf{x}=\mathbf{y}\}$ is maximal over all couplings, and is equal to one minus
$|P(\mathbf{x},\cdot)-P(\mathbf{y},\cdot)|_{\text{TV}}=\int\min\left\{\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})},\;\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{y})},\;1\right\}\bar{q}(\mathrm{d}\mathbf{x}^{\star}).$ (16)
  1. 1.
    Draw $\mathbf{x}^{\star}\sim\bar{q}$.
  2. 2.
    Draw $u$ from a Uniform(0,1) distribution.
  3. 3.
    If $u<{\hat{Z}(\mathbf{x}^{\star})}/{\hat{Z}(\mathbf{x})}$, set $\mathbf{x}^{\prime}=\mathbf{x}^{\star}$, otherwise set $\mathbf{x}^{\prime}=\mathbf{x}$.
  4. 4.
    If $u<{\hat{Z}(\mathbf{x}^{\star})}/{\hat{Z}(\mathbf{y})}$, set $\mathbf{y}^{\prime}=\mathbf{x}^{\star}$, otherwise set $\mathbf{y}^{\prime}=\mathbf{y}$.
  5. 5.
    Return $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$
Algorithm 4 Common draws coupling of IMH, denoted by $\bar{P}$, for chains currently at $(\mathbf{x},\mathbf{y})$.
Let $(\mathbf{x}_{t},\mathbf{y}_{t})$ be a coupled chain started from $(\mathbf{x},\mathbf{y})$ and evolving according to $\bar{P}$. Denoting the meeting time by
$\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t}\},$ (17)
the coupling inequality states that, for $t\geq 1$,
$|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\leq\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t),$ (18)
where the probability $\mathbb{P}_{\mathbf{x},\mathbf{y}}$ is under the law of $(\mathbf{x}_{t},\mathbf{y}_{t})$ started from $(\mathbf{x},\mathbf{y})$ at time zero. We will relate the probability $\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)$ to the rejection probabilities of IMH from $\mathbf{x}$ and $\mathbf{y}$, and we define
$r:\mathbf{x}\mapsto\int_{\mathbf{x}^{\star}\neq\mathbf{x}}\left(1-\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})\right)\bar{q}(\mathrm{d}\mathbf{x}^{\star}),$ (19)
where $\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})$ is defined in (10).
The meeting time $\tau$ is the first time at which both chains accept the proposal simultaneously, which corresponds to the first time at which the chain with the highest weight accepts the proposal. Indeed, if $\hat{Z}(\mathbf{x})\geq\hat{Z}(\mathbf{y})$, then $\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})\leq\alpha_{\text{RH}}(\mathbf{y},\mathbf{x}^{\star})$ for all $\mathbf{x}^{\star}$, and thus $u<\alpha_{\text{RH}}(\mathbf{x},\mathbf{x}^{\star})$ implies that $u<\alpha_{\text{RH}}(\mathbf{y},\mathbf{x}^{\star})$. Thus, conditionally on $\mathbf{x}_{0}=\mathbf{x},\mathbf{y}_{0}=\mathbf{y}$, the meeting time $\tau$ follows a Geometric distribution with parameter $1-r(\mathbf{x})$, where $r(\mathbf{x})$ is defined in (19). Recall that the survival function of a Geometric variable $T$ with parameter $\gamma$ is given by: $\mathbb{P}(T>t)=(1-\gamma)^{k}$ for $t\in\mathbb{N}$. Still assuming $\hat{Z}(\mathbf{x})\geq\hat{Z}(\mathbf{y})$, we obtain, for $t\geq 1$,
$|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\leq\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)=(r(\mathbf{x}))^{t}.$ (20)
The above upper bound is given in Roberts & Rosenthal (2011). In their remark following Theorem 5, they state that this is also a lower bound without providing a proof. We do so below, for both discrete and continuous state spaces; Roberts & Rosenthal (2011) focus on non-atomic spaces. First, we express
$|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}=\sup_{A\in\mathcal{B}(\mathbb{R}^{d})}|P^{t}(\mathbf{x},A)-P^{t}(\mathbf{y},A)|,$ (21)
and we select the set $A=\mathbb{R}^{d}\setminus\{\mathbf{x}\}$ to obtain a lower bound. Then $P^{t}(\mathbf{y},A)$ = 1 since $\mathbf{x}\neq\mathbf{y}$ and assuming that $q(\{\mathbf{x}\})=0$, while $P^{t}(\mathbf{x},A)=1-\left(r(\mathbf{x})\right)^{t}$, i.e. the chain is in $A$ at step $t$ except if $t$ proposals have been rejected. The situation is slightly more complicated if the proposal has non-zero mass on $\{\mathbf{x}\}$ and $\{\mathbf{y}\}$, i.e. in discrete state spaces, but the following result still holds. The proof is in Appendix A.3.
Theorem 3.1.
Let $(\mathbf{x}_{t},\mathbf{y}_{t})$ be a Markov chain evolving according to $\bar{P}$ in Algorithm 4 and starting from $\mathbf{x}_{0}=\mathbf{x}$ and $\mathbf{y}_{0}=\mathbf{y}$. Let $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t}\}$, and let $r(\mathbf{x})$ be defined as in (19). Then, for all $t\geq 1$,
$|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}=\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)=\max(r(\mathbf{x}),r(\mathbf{y}))^{t}.$ (22)
Thus, the chain $(\mathbf{x}_{t},\mathbf{y}_{t})$ generated by the common draws coupling follows a maximal coupling, as in Pitman (1976): the coupling inequality is an equality. To the best of our knowledge, this is the only known case of “all time maximal” couplings of an MCMC algorithm. Note also that the upper bound in (22) decreases geometrically in $t$. The polynomial rates come later, when we integrate over $\mathbf{x}$ or $\mathbf{y}$.
4 Meeting times and polynomial convergence
4.1 Meeting times of lagged chains
We consider coupled IMH chains with a lag, as in Middleton et al. (2019). The construction is described in Algorithm 5. The generated chains $(\mathbf{x}_{t})_{t\geq 0}$ and $(\mathbf{y}_{t})_{t\geq 0}$ have the same marginal distribution, that of an IMH chain started from $\bar{q}$.
  1. 1.
    Set $\tau=+\infty$ and $t=1$.
  2. 2.
    Draw $\mathbf{x}_{0}\sim\bar{q}$ and $\mathbf{y}_{0}\sim\bar{q}$ independently.
  3. 3.
    Draw $u$ from a $\text{Uniform}(0,1)$ distribution.
  4. 4.
    If $u<\hat{Z}(\mathbf{y}_{0})/\hat{Z}(\mathbf{x}_{0})$, set $\mathbf{x}_{1}=\mathbf{y}_{0},\tau=1$. Otherwise, set $\mathbf{x}_{1}=\mathbf{x}_{0}$.
  5. 5.
    While $\tau=+\infty$
    1. (a)
      Set $t=t+1$.
    2. (b)
      Sample $(\mathbf{x}_{1+t},\mathbf{y}_{t})\sim\bar{P}((\mathbf{x}_{t},\mathbf{y}_{t-1}),\cdot)$, the common draws coupling of PIMH in Algorithm 4.
    3. (c)
      If $\mathbf{x}_{1+t}=\mathbf{y}_{t}$, set $\tau=1+t$.
  6. 6.
    Return $\tau,\mathbf{x}_{0},\mathbf{y}_{0},\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{y}_{\tau-1},\mathbf{x}_{\tau}$.
Algorithm 5 Coupled PIMH with a lag.
We relate the distribution of the meeting times generated by Algorithm 5 to the expected rejection probability in the following result.
Proposition 4.1.
Consider $\tau$ generated by Algorithm 5. Under Assumption 1, for all $t\geq 1$, we have
$\mathbb{P}(\tau>t)\leq\mathbb{E}_{\bar{q}}\left[\left(r(\mathbf{x})\right)^{t}\right].$ (23)
This connection between meeting times and expected rejection probability motivates our subsequent analysis of the expected rejection probability, which appears central in the study of IMH (e.g. Theorem 6 in Roberts & Rosenthal (2011)). Our bounds are explicit functions of $t$ and $N$.
Proposition 4.2.
Fix $p\geq 2$ and let
$\beta_{p}:=1-\frac{1}{2^{\frac{3p-2}{p-1}}q(\omega^{p})^{\frac{1}{p-1}}}.$ (24)
Under Assumptions 1-2, there exist constants $A_{p},C_{p}>0$, depending only on $p$ and $q(\omega^{p})$, such that for all $N\geq 1$, for all $t\geq 1$, the following holds:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\right]\leq\frac{A_{p}}{N^{(t\wedge p)/2}}\beta_{p}^{t}+\frac{C_{p}}{t^{p}N^{p/2}}.$ (25)
Proposition 4.2 holds for all $t\geq 1$ and all $N\geq 1$. The bounds decay to 00 as either $N$ or $t$ approaches infinity, polynomially with rate at most $N^{-1/2}$ w.r.t. $N$, and, for fixed $N$, polynomially with rate $t^{-p}$ where $p$ is the number of assumed moments of $\omega$ under $q$. A direct consequence of the previous two propositions is a bound on the tails of the meeting times.
Proposition 4.3.
Consider $\tau$ generated by Algorithm 5. Under Assumptions 1-2, there exists $C$ such that for all $N\geq 1$ and all $t\geq 1$, if $p\geq 2$ in Assumption 2,
$\mathbb{P}(\tau>t)\leq\frac{C}{\sqrt{N}t^{p}}.$ (26)
That bound retains the slowest rates in $N$ and $t$ from the previous result. Proposition 4.3 is consistent with Proposition 8 in Middleton et al. (2019), which showed that $\mathbb{P}(\tau=1)$ approaches one as $N\to\infty$ under the assumption of bounded weights. However, our present assumptions are considerably weaker, and we provide explicit dependencies on both $N$ and $t$.
Remark 4.1.
We comment on the sharpness of the dependency on $N$ in Proposition 4.3. For $t=1$, the result reads $\mathbb{P}(\tau>1)\leq C/\sqrt{N}$. The event $\{\tau>1\}$ corresponds to the rejection of $\mathbf{x}^{\star}$ from a state $\mathbf{x}$, both $\mathbf{x},\mathbf{x}^{\star}$ being independent draws from $\bar{q}$. Here we show that we cannot improve upon the rate $N^{-1/2}$ as a function of $N$. The central limit theorem implies $\sqrt{N}(\hat{Z}_{N}(\mathbf{x})-1)\to\text{Normal}(0,q(\omega^{2})-1)$ in distribution. Therefore, $\mathbb{P}(\hat{Z}_{N}(\mathbf{x})\geq 1+N^{-1/2})\rightarrow p_{0}$ as $N\to\infty$, with $p_{0}$ depending on $q(\omega^{2})$. The same argument shows $\mathbb{P}(\hat{Z}_{N}(\mathbf{x}^{\star})\leq 1-N^{-1/2})\rightarrow p_{1}$ as $N\to\infty$, with $p_{1}$ depending on $q(\omega^{2})$. Therefore, we can choose a large enough $N$ that depends on $q(\omega^{2})$ such that $\mathbb{P}(\hat{Z}_{N}(\mathbf{x})\geq 1+N^{-1/2})\geq p_{0}/2$ and $\mathbb{P}(\hat{Z}_{N}(\mathbf{x}^{\star})\leq 1-N^{-1/2})\geq p_{1}/2$. Thus, with a constant probability $c$, $Z_{N}(\mathbf{x}^{\star})\leq 1-N^{-1/2}$ and $\hat{Z}_{N}(\mathbf{x})\geq 1+N^{-1/2}$ occur simultaneously, and thus the acceptance probability is at most $(1-N^{-1/2})/(1+N^{1/2})\leq 1-N^{-1/2}$. In turn this means that the rejection probability is at least $cN^{-1/2}$.
4.2 Polynomial convergence rates
As discussed in Section 6 of Jacob et al. (2020) and in Biswas et al. (2019), lagged chains such as those generated by Algorithm 5 can be employed to bound the total variation distance between the chain at time $t$ and its stationary distribution. We aim for bounds on $\left|P^{t}(\mathbf{x},\cdot)-\pi\right|_{\mathrm{TV}}$ that are explicit in their dependency on the iteration $t$ and the number of particles $N$. We present the following result.
Theorem 4.1.
Consider $\tau$ generated by Algorithm 5. Let $P$ be the transition kernel of the PIMH chain as in Algorithm 3. Under Assumption 1, we have that for all $t\geq 0$,
$\left|\bar{q}P^{t}-\bar{\pi}\right|_{\text{TV}}\leq\mathbb{E}\left[\max\left(0,\tau-1-t\right)\right].$ (27)
Furthermore, if Assumption 2 also holds, there exists a constant $C$, independent of $t$ and $N$, such that for all $N\geq 1$ and $t\geq 0$,
$\left|\bar{q}P^{t}-\bar{\pi}\right|_{\text{TV}}\leq\frac{C}{\sqrt{N}(1+t)^{p-1}}.$ (28)
Remark 4.2.
The case $t=0$ states that $|\bar{q}-\bar{\pi}|_{\text{TV}}\leq CN^{-1/2}$, which may seem strange as both $\bar{\pi}$ and $\bar{q}$ defined in (7)-(8) are defined on spaces growing with $N$. With the density representation of the total variation distance and $\bar{\pi}(\mathbf{x})=\hat{Z}(\mathbf{x})\bar{q}(\mathbf{x})$, we can directly compute
$|\bar{q}-\bar{\pi}|_{\text{TV}}$ $=\frac{1}{2}\int|1-\hat{Z}(\mathbf{x})|\bar{q}(\mathrm{d}\mathbf{x})=\frac{1}{2}\mathbb{E}_{\bar{q}}\left[|1-\hat{Z}(\mathbf{x})|\right]\leq\frac{1}{2}\mathbb{E}_{\bar{q}}\left[|1-\hat{Z}(\mathbf{x})|^{2}\right]^{1/2}\leq\frac{1}{2}M(2)^{1/2}N^{-1/2},$ (29)
where the first inequality is Cauchy–Schwarz and the second uses Proposition 1.1 under Assumption 2. Furthermore, in the large $N$ asymptotics we expect $1-\hat{Z}(\mathbf{x})$ to behave as a Normal distribution with mean zero and standard deviation $\sqrt{q(\omega^{2})-1}/\sqrt{N}$, so that its expectation should indeed behave as $\sqrt{(2/\pi)(q(\omega^{2})-1)}/\sqrt{N}$.
We can also state a bound for the convergence of the chain started at any initial point $\mathbf{x}\in\mathbb{X}$.
Corollary 4.1.
Under Assumptions 1-2, there exists a constant $\tilde{C}$, independent of $t$ and $N$, such that for all $N\geq 1$, $t\geq 1$, and any starting point $\mathbf{x}\in\mathbb{X}^{N}$,
$\left|P^{t}(\mathbf{x},\cdot)-\bar{\pi}\right|_{\text{TV}}\leq(r(\mathbf{x}))^{t}+\frac{\tilde{C}}{\sqrt{N}(1+t)^{p-1}}.$ (30)
Theorem 4.1 and Corollary 4.1 provide explicit bounds on the convergence rate of the PIMH algorithm. Both results are interpretable in terms of the number of iterations $t$ and the number of particles $N$, and apply to IMH as a special case when $N=1$. The difference between these results lies in the starting distribution. Practitioners would typically start the algorithm from the proposal distribution, as it is the best available approximation of the target. Corollary 4.1 reveals two phases in the convergence: an initial phase where the distance decays exponentially in $t$ but not arbitrarily with $N$, followed by a polynomial decay in both $t$ and $N$.
We add a result for the case $N=1$ i.e. standard IMH, which holds under the assumption that $q(\omega^{p})<\infty$ for $p>1$, whereas Corollary 4.1 requires $p\geq 2$ in Assumption 2. The proof is in Appendix A.4.3. As discussed in Section 4.3 the result in Proposition 4.4 is similar to existing results in the literature, although we have not found statements expressed as simply, and our proof appears to be original.
Proposition 4.4.
Consider IMH under Assumption 1, and assume $q(\omega^{p})<\infty$ for $p>1$. There exists a constant $D$ independent of $t$ such that for all $t\geq 1$, and any starting point $x\in\mathbb{X}$,
$\left|P^{t}(x,\cdot)-\pi\right|_{\text{TV}}\leq(r(x))^{t}+\frac{D}{(1+t)^{p-1}}.$ (31)
The purpose of the following example is to demonstrate that the rate $t^{-(p-1)}$ in Corollary 4.1 and Proposition 4.4 cannot be improved beyond polylogarithmic factors. The proof is provided in Appendix A.4.4.
Example 3.
Consider the IMH algorithm targeting $\pi(x):=Z_{\pi}x^{-p}$ on $[2,\infty)$, with proposal distribution $q(x):=Z_{q}\log^{2}(x)/x^{-(p+1)}$ on $[2,\infty)$, started from $x_{0}=3$. If $p\geq 2$, Assumption 2 holds with that $p$, and there exist $C<\infty$ and $t_{0}\in\mathbb{N}$ such that, for all $t\geq t_{0}$,
$\left|P^{t}(x_{0},\cdot)-\pi\right|_{\text{TV}}$ $\geq\frac{C}{t^{p-1}(\log t)^{3(p-1)}}.$
4.3 Related results on IMH
The convergence of IMH has garnered significant interest over decades, and in particular the sub-geometric rates have been studied in several works including Jarner & Roberts (2002); Douc et al. (2007); Roberts & Rosenthal (2011); Andrieu et al. (2022). One approach utilizes drift and minorization techniques (Jarner & Roberts, 2002).
Theorem 4.2 (Theorem 5.3 in Jarner & Roberts (2002)).
Let $P$ be the transition kernel of the IMH chain as in Algorithm 2. Assume that for some $r>0$,
$\displaystyle\pi(A_{\epsilon})=\mathcal{O}\left(\epsilon^{1/r}\right)\quad\text{for}\quad\epsilon\to 0,$ (32)
where $A_{\epsilon}=\left\{x\in\mathbb{X}:\omega(x)>1/\epsilon\right\}$, for any $\epsilon>0$. Then, for any $x\in\mathbb{X}$, and any $t\geq 1$, we have that
$\displaystyle\lim_{t\to\infty}(1+t)^{\beta}\left|P^{t}(x,\cdot)-\pi\right|_{\mathrm{TV}}=0,$ (33)
for any $0\leq\beta\leq\frac{s-r}{r}$, with $r<s<r+1$.
The $\mathcal{O}$ notation here is such that if $f(x)=\mathcal{O}(g(x))$ then there exists a constant $M$ such that $|f(x)|\leq M|g(x)|$ for all $x$ in the domain of $f$. Theorem 4.2 provides a polynomial rate of convergence for the IMH chain in total variation of order $o\left(t^{-1/r+\kappa}\right)$ for any $\kappa>0$ under the assumption that the tail weights satisfy the condition specified in equation (32). Notably, under the assumption $q(\omega^{p})<\infty$, the condition in (32) is satisfied for $r=1/(p-1)$, using Markov’s inequality. Our Proposition 4.4 differs slightly as our bounds are in $t^{-(p-1)}$ instead of $t^{-(p-1)+\kappa}$ for some arbitrarily small $\kappa>0$. Similar results can be obtained using weak Poincaré inequalities as described in Remark 29 of Andrieu et al. (2022), under $\pi(\omega^{p})$ with $p>1$ which amounts to our Assumption 2 with $p>2$.
Our bounds in Theorem 4.1 and Corollary 4.1 have the advantage of providing an explicit dependency on $N$ in the case of PIMH, which is critical for the results on bias removal in Section 5.
4.4 Comparing the biases of IS and IMH
We can now compare the bias of IMH and IS under the assumption $q(\omega^{p})<\infty$ for $p\geq 2$. Suppose that the goal is to obtain a single sample close to $\pi$ in total variation distance, with a budget of $N$ samples from $q$ and $N$ evaluations of $\omega$.
  • One approach is sampling-importance resampling (SIR), which refers to the following procedure. First, run Algorithm 1. Then, draw $k\sim\text{Categorical}(\omega(x_{1}),\ldots,\omega(x_{N}))$ and return $x_{k}$. For a test function $f$ with $|f|_{\infty}\leq 1$, under the conditions of Theorem 2.1 the marginal distribution $\mu_{N}^{\text{SIR}}$ of $x_{k}$ satifies
    $\mu_{N}^{\text{SIR}}(f)-\pi(f)\sim_{N\to\infty}-q(\omega^{2}\cdot(f-\pi(f)))N^{-1}.$ (34)
    Skare et al. (2003) provide a similar result in the case of bounded weights, and propose a modification of SIR to reduce this bias to $N^{-2}$.
  • On the other hand, Theorem 4.1 suggests that IMH (with one particle) after $N$ iterations provides a sample from a distribution $qP^{N-1}$, for which, under the condition $q(\omega^{p})<\infty$,
    $\sup_{f:|f|_{\infty}\leq 1}\left\{qP^{N-1}(f)-\pi(f)\right\}\leq CN^{-(p-1)}.$ (35)
    Thus, the bias may be reduced faster when increasing the computing budget in IMH rather than in SIR, as soon as $p>2$; for the modified SIR of Skare et al. (2003), IMH is faster as soon as $p>3$.
In the case of bounded weights, MCMC methods such as PIMH or particle Gibbs (Andrieu et al., 2010) are geometrically ergodic (e.g. Lee et al., 2020) and the bias comparison is clearly at the advantage of MCMC algorithms, as discussed in Cardoso et al. (2022).
5 Bias removal for self-normalized importance sampling
5.1 Construction
The bias of importance sampling was described in Section 2, and that of IMH in Section 4. Here we consider the removal of the bias, and the associated cost. For this purpose, Middleton et al. (2019) employ common random numbers couplings of PIMH and the approach of Glynn & Rhee (2014). We pursue this strategy.
Upon running Algorithm 5 with $N\geq 1$, with $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t-1}\}$, one can compute the following unbiased estimator:
$\hat{F}_{u}=\hat{F}(\mathbf{x}_{0})+\sum_{t=1}^{\tau-1}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\},$ (36)
where $\hat{F}:\mathbf{x}\mapsto\sum_{n=1}^{N}\omega(x_{n})f(x_{n})/(\sum_{n=1}^{N}\omega(x_{n}))$ and $f$ is a test function. By convention the sum in (36) is zero in the event $\{\tau=1\}$, and it is also equal to the infinite sum $\sum_{t=1}^{\infty}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\}$ since $\hat{F}(\mathbf{x}_{t})=\hat{F}(\mathbf{y}_{t-1})$ from time $\tau$ onward. The lack of bias can be seen via a telescopic sum argument, since $\mathbf{x}_{t}$ and $\mathbf{y}_{t}$ have the same marginal distribution for all $t$, and provided that limit and expectation can be swapped. Since $\mathbf{x}_{0}\sim\bar{q}$, $\hat{F}(\mathbf{x}_{0})$ is the (biased) IS estimator. In contrast, $\hat{F}_{u}$ in (36) is unbiased, under some conditions. Middleton et al. (2019) consider the case where $\omega$ is uniformly upper bounded, and they show that (36) can have a finite variance. Below we work under the weaker Assumptions 1-2, and we derive results on the moments of unbiased IS and on its comparison with regular IS.
Remark 5.1.
(36) is an instance of unbiased MCMC (Jacob et al., 2020; Atchadé & Jacob, 2024), where here MCMC is PIMH, and various generic improvements could be considered, such as increasing the lag between the chains, or introducing a burn-in parameter. However, in the particular case of PIMH, the number of particles $N$ is a key parameter and here we focus on the regime $N\to\infty$, in which case $\hat{F}_{u}$ naturally compares with $\hat{F}(\mathbf{x}_{0})$, which is the regular IS estimator. Hence we call (36) the unbiased self-normalized importance sampling estimator (UIS), but it could also be called unbiased PIMH.
5.2 Moments of unbiased self-normalized importance sampling
We subtract $\pi(f)$ from all terms in (36) to obtain
$\hat{F}_{u}-\pi(f)=\hat{F}(\mathbf{x}_{0})-\pi(f)+\sum_{t=1}^{\infty}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\}\mathds{1}(\tau>t).$ (37)
We introduce the notation
$\Delta_{t}$ $=\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1}),\quad\text{BC}=\sum_{t=1}^{\infty}\Delta_{t}\mathds{1}(\tau>t),$ (38)
where BC stands for the bias cancellation term. Using Minkowski’s inequality, the moments of the error of $\hat{F}_{u}$ can be bounded by the moments of the error of the IS estimator $\hat{F}(\mathbf{x}_{0})$, as in Theorem 2.3, and the moments of BC.
A first result is that, for bounded test functions $f$, $\hat{F}_{u}$ has as many moments as the meeting time $\tau$, which is nearly $p$ under Assumption 2. The proof is in Appendix A.5.1.
Proposition 5.1.
Assume that $|f|_{\infty}\leq 1$ and let $s\geq 1$. If Assumptions 1-2 hold with $p\geq 2$ and $p>s$, then the meeting time $\tau$ has $s$ finite moments, and the unbiased self-normalized importance sampling (UIS) estimator $\hat{F}_{u}$ in (36) has $s$ finite moments.
To deal with unbounded test functions, we need to control the moments of the terms $\Delta_{t}$. For this we derive the following result about PIMH at any iteration $t$, under the same conditions as Theorem 2.3. The proof is in Appendix A.5.2.
Proposition 5.2.
Assume that there exist $p\in[2,\infty)$ and $r\in[2,\infty]$ such that $q(\omega^{p})<\infty$ and $q(f^{r})<\infty$, and $q(f^{2}\cdot\omega^{2})<\infty$. Let $(\mathbf{x}_{t})$ be the PIMH chain started from $\mathbf{x}_{0}\sim\bar{q}$. Then, for any $2\leq s\leq pr/(p+r+2)$ and any $N\geq 1$, there exists $C$ such that for all $t\geq 0$:
$\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[\left\lvert\hat{F}(\mathbf{x}_{t})-\pi(f)\right\rvert^{s}\right]\leq CN^{-s/2},$
where the constant $C$ depends on $r,p,s,q(f^{r}),q(\omega^{p}),q(f^{2}\cdot\omega^{2})$. When $r=\infty$, the statement holds for $f$ such that $|f|_{\infty}<\infty$ and all $s\leq p$.
By Minkowski’s inequality, under the conditions of Proposition 5.2, the moments of $\Delta_{t}$ have similar bounds. This can be used to obtain the following result, proven in Appendix A.5.3.
Proposition 5.3.
Assume that there exist $p\in(2,\infty)$ and $r\in[2,\infty]$ such that $q(\omega^{p})<\infty$ and $q(|f|^{r})<\infty$, and $q(f^{2}\cdot\omega^{2})<\infty$, then for any $2\leq s<p$ such that $pr/(p+r+2)>ps/(p-s)$, and for any $N\geq 1$, the unbiased importance sampling (UIS) estimator satisfies:
$\mathbb{E}\left[\left\lvert\hat{F}_{u}-\pi(f)\right\rvert^{s}\right]\leq CN^{-s/2},$
where the constant $C$ depends on $r,p,s,q(|f|^{r}),q(\omega^{p}),q(f^{2}\cdot\omega^{2})$. When $r=\infty$, the statement holds for $f$ such that $|f|_{\infty}<\infty$ and all $s<p$ such that $p>sp/(p-s)$.
Two finite moments ($s=2$ in the above results) are particularly useful. Indeed, a confidence interval for $\pi(f)$ can be constructed from independent copies of $\hat{F}_{u}$ using the Central Limit Theorem as long as $\mathbb{V}[\hat{F}_{u}]<\infty$. Beyond the empirical average, robust mean estimation strategies could be envisioned. The typical assumption in robust mean estimation is that users have access to a random variable $X$ with mean $\mu$ of interest, and a finite variance $\sigma^{2}$ that may be unknown. The goal is to aggregate independent copies of $X$ into an estimator of $\mu$ with lighter tails than the empirical average. Such results can be found in Devroye et al. (2016); Lugosi & Mendelson (2019b, a); Lecué & Lerasle (2020); Minsker & Ndaoud (2021). These techniques are directly applicable with $X=\hat{F}_{u}$ and $\mu=\pi(f)$, whereas the application of robust mean estimation to self-normalized importance sampling is not a priori straightforward, and was explored in Dau (2022) in the case of median-of-means.
5.3 The asymptotic price of bias removal
We consider the price of debiasing self-normalized importance sampling in terms of inefficiency, defined by the variance multiplied by the average cost. We start by comparing the mean squared errors of unbiased and regular IS. From (37), we take the square and use Cauchy–Schwarz to obtain
$\left|\mathbb{E}\left[(\hat{F}_{u}-\pi(f))^{2}\right]-\mathbb{E}\left[(\hat{F}(\mathbf{x}_{0})-\pi(f))^{2}\right]\right|$ $\leq\sqrt{\mathbb{E}\left[(\hat{F}(\mathbf{x}_{0})-\pi(f))^{2}\right]\cdot\mathbb{E}[\text{BC}^{2}]}+\mathbb{E}[\text{BC}^{2}],$ (39)
with $\text{BC}=\sum_{t=1}^{\tau-1}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\}$.
The mean squared error (MSE) of IS, which is the term $\mathbb{E}[(\hat{F}(\mathbf{x}_{0})-\pi(f))^{2}]$, is of order $N^{-1}$ under conditions stated in Theorem 2.3. If we can bound $\mathbb{E}[\text{BC}^{2}]$ by a term that decreases faster than $N^{-1}$, then the MSE of UIS would be asymptotically equivalent to that of IS. Intuitively, the bias cancellation term goes to zero for two reasons: first because $\tau$ goes to one as $N\to\infty$, and the bias cancellation term equals zero in the event $\{\tau=1\}$. Secondly, each term $\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})$ goes to 00 as $N\to\infty$, under the conditions of Proposition 5.2. We obtain the following result, proven in Appendix A.5.4.
Proposition 5.4.
Let $\hat{F}_{u}$ be the UIS estimator defined as (36) and $\hat{F}(\mathbf{x})$ with $\mathbf{x}\sim\bar{q}$ be the IS estimator. Suppose that the assumptions of Proposition 5.3 are satisfied with $s=2$, that is: $p>2$ and $r>2$ such that $q(\omega^{p})<\infty$ and $q(|f|^{r})<\infty$, and $q(f^{2}\cdot\omega^{2})<\infty$, with $2p+4r+4<r\cdot p$. Then the mean squared error of $\hat{F}_{u}$ and that of $\hat{F}(\mathbf{x})$ are asymptotically equivalent:
$\lim_{N\to\infty}N\cdot\mathbb{E}\left[(\hat{F}_{u}-\pi(f))^{2}\right]=\lim_{N\to\infty}N\cdot\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[(\hat{F}(\mathbf{x})-\pi(f))^{2}\right].$
The assumption $2p+4r+4<r\cdot p$ is for example satisfied if $r=\infty$ and $p=4+\epsilon$ with an arbitrary $\epsilon>0$, or if $p=5$ and $r=15$. However it cannot be satisfied with $p\leq 4$.
The cost of UIS is that of running Algorithm 5. It starts with two draws from $\bar{q}$, i.e. $2N$ draws from $q$, and as many evaluations of the weight function $\omega$. Then either $\tau=1$ or the algorithm enters its while loop up to the meeting time $\tau$, drawing $N$ new particles at each iterate of the loop. Counting the cost in terms of the number of evaluations of $\pi$, UIS has an overall cost of $\mathcal{C}=2N+N(\tau-1)$. If Assumption 2 holds with $p\geq 2$, using Proposition 4.3 then $\mathbb{E}[\tau]=\sum_{t\geq 0}\mathbb{P}(\tau>t)$ is finite, and furthermore, $\mathbb{E}[\mathcal{C}]\sim_{N\to\infty}2N$. We summarise these first observations in the next statement.
Proposition 5.5.
The cost of the UIS estimator $\hat{F}_{u}$ in (36) is $\mathcal{C}=2N+N(\tau-1)$, and, under Assumption 2 for $p\geq 2$, there exists a constant $C$ such that $\mathbb{E}[\tau]\leq 1+C/\sqrt{N}$.
Combining Proposition 5.4 with the above result, we see that under some conditions, the inefficiency $\mathbb{E}[\mathcal{C}]\mathbb{V}[\hat{F}_{u}]$ is equivalent to twice that of self-normalized importance sampling as $N\to\infty$. Indeed, the mean squared errors are equivalent but the cost of UIS behaves as $2N$ instead of $N$ for IS. In Algorithm 5 the $N$ particles in $\mathbf{y}_{0}$ are required to determine $\tau$ but in the event $\{\tau=1\}$ they do not participate directly in the estimator $\hat{F}_{u}$.
5.4 An improved unbiased estimator
A trick provides a remedy, and cuts the asymptotic inefficiency by a half. We can view $\hat{F}_{u}$ as a deterministic function of initial states $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ drawn from $\bar{q}$ independently, as well as extra variables (a sequence of proposals from $\bar{q}$, a sequence of uniform random variables) that we collectively label $\zeta$; we denote $\hat{F}_{u}$ by $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$. Then we define the symmetrized UIS estimator:
$\tilde{F}_{u}=\frac{1}{2}\left(\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)+\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)\right).$ (40)
Computing (40) only requires simple modifications of Algorithm 5. Indeed, either $\hat{Z}(\mathbf{x}_{0})\geq\hat{Z}(\mathbf{y}_{0})$ or $\hat{Z}(\mathbf{x}_{0})<\hat{Z}(\mathbf{y}_{0})$. In the first case, we always have $\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)=\hat{F}(\mathbf{y}_{0})$, and $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$ can be computed following Algorithm 5 and (36). In the second case, we always have $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)=\hat{F}(\mathbf{x}_{0})$, and $\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)$ can be computed following Algorithm 5 and (36) with the role of $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ swapped. That trick amounts to a Rao–Blackwellization over the arbitrary specification of which draws from $\bar{q}$ are used as $\mathbf{x}_{0}$ and as $\mathbf{y}_{0}$. The following statement is a mild variation of the previous results and is stated without a proof.
Proposition 5.6.
Consider the symmetrized UIS estimator $\tilde{F}_{u}$ in (40). Suppose that the conditions of Proposition 5.4 are satisfied. Then $\tilde{F}_{u}$ is an unbiased estimator of $\pi(f)$ for any $N\geq 1$, with finite expected cost and finite variance, and its inefficiency is equivalent to that of IS as $N\to\infty$.
The result supports the intuition that $\tilde{F}_{u}$ should be preferred to $\hat{F}_{u}$ in practice, and that the asymptotic efficiency of $\tilde{F}_{u}$ matches that of self-normalized importance sampling as $N\to\infty$, under some conditions on $\omega$ and $f$ such as those in Proposition 5.4.
Acknowledgements.
Through the second author and CY Initiative Emergence, this project was supported by France 2030.
Appendix A Proofs
A.1 Proofs of Section 1
A.1.1 Proof of (11)
We introduce the following extended state space justification of PIMH. Let the target and proposal distributions be
$\tilde{\pi}(x_{1},\ldots,x_{N},k)$ $=\frac{1}{N}\pi(x_{k})\prod_{n\neq k}q(x_{n}),$ (41)
$\tilde{q}(x_{1},\ldots,x_{N},k)$ $=\frac{\omega(x_{k})}{\sum_{n=1}^{N}\omega(x_{n})}\prod_{n=1}^{N}q(x_{n}).$ (42)
The state space is $\mathbb{X}^{N}\times\{1,\ldots,N\}$, i.e. $k$ is an integer in $\{1,\ldots,N\}$. The ratio of target divided by proposal on the extended space takes the following form:
$\tilde{\omega}(x_{1},\ldots,x_{N},k)=\frac{\tilde{\pi}(x_{1},\ldots,x_{N},k)}{\tilde{q}(x_{1},\ldots,x_{N},k)}=\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n}).$ (43)
In other words, the ratio simplifies to $\hat{Z}(\mathbf{x})$, thus IMH with proposal $\tilde{q}$ and target $\tilde{q}$ is equivalent to Algorithm 3.
Draws from the proposal $\tilde{q}(x_{1},\ldots,x_{N},k)$ can be obtained by first sampling $x_{1},\ldots,x_{N}$ i.i.d. from $q$, then computing the importance weights $\omega(x_{n})$ for $n=1,\ldots,N$, and finally sampling $k$ from a Categorical distribution with probabilities given by $\omega(x_{n})/\sum_{m=1}^{M}\omega(x_{m})$. Under the extended target $\tilde{\pi}$, we have the following marginals and conditionals.
  • The marginal distribution of $k$ is the Uniform distribution on $\{1,\ldots,N\}$.
  • The marginal distribution of $x_{1},\ldots,x_{N}$ is obtained as
    $\tilde{\pi}(x_{1},\ldots,x_{N})=\sum_{n=1}^{N}\frac{1}{N}\pi(x_{n})\prod_{m\neq n}q(x_{m}).$ (44)
  • The conditional distribution of $k$ given $x_{1},\ldots,x_{N}$ can be obtained by dividing (41) by (44),
    $\tilde{\pi}(k|x_{1},\ldots,x_{N})=\frac{\omega(x_{k})}{\sum_{n=1}^{N}\omega(x_{n})}.$ (45)
  • The conditional distribution of $x_{1},\ldots,x_{N}$ given $k$ can be obtained by dividing (41) by $N^{-1}$, and is given by
    $\tilde{\pi}(x_{1},\ldots,x_{N}|k)=\pi(x_{k})\prod_{n\neq k}q(x_{n}),$ (46)
    and in particular $\tilde{\pi}(x_{k}|k)$ is $\pi(x_{k})$.
Note that $f(x_{k})$ can be written as a more explicit function of $(x_{1},\ldots,x_{N},k)$, as $\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)$. We use the above properties to obtain the following equalities, as in Andrieu et al. (2010), Appendix B.5, either conditioning on $k$ in the first line, or on $x_{1},\ldots,x_{N}$ in the second line:
$\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)]$ $=\mathbb{E}_{\bar{\pi}}[\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)|k]]=\mathbb{E}_{\pi}[f],$ (47)
$=\mathbb{E}_{\bar{\pi}}[\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)|x_{1},\ldots,x_{N}]]=\mathbb{E}_{\bar{\pi}}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}\right].$ (48)
We obtain (11), i.e. the equality $\mathbb{E}_{\pi}[f]=\mathbb{E}_{\mathbf{x}\sim\bar{\pi}}[\hat{F}(\mathbf{x})]$, with $\hat{F}$ as in (4).
A.1.2 Proofs of Proposition 1.1 and Lemma 1.1
Proof of Proposition 1.1.
Using the result of Petrov (1975), Section III.5 (item 16, p. 60): for $X_{1},\ldots,X_{N}$ independent variables with zero mean and $p$ finite moments, $p\geq 2$, we have
$\mathbb{E}\left[\left|\sum_{n=1}^{N}X_{n}\right|^{p}\right]\leq m(p)N^{p/2-1}\sum_{n=1}^{N}\mathbb{E}[|X_{n}|^{p}],$
where $m(p)$ is a positive number depending only on $p$. As described in Ren & Liang (2001), the constant $m(p)$ satisfies $(m(p))^{1/p}\leq p-1$; in fact they provide a sharper bound, but we do not need it here. For i.i.d. variables the right-hand side becomes $m(p)N^{p/2}\mathbb{E}[|X_{1}|^{p}]$. If we consider the average instead of the sum on the left, then the right-hand side becomes $m(p)N^{-p/2}\mathbb{E}[|X_{1}|^{p}]$. Since $q(\omega)=1$ and assuming that $q(\omega^{p})<\infty$, we define $X_{n}=\omega_{n}-1$ and apply the above result to obtain
$\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1|^{p}\right]\leq(p-1)^{p}N^{-p/2}q((\omega-1)^{p}).$
Next we can use the $C_{p}$-inequality, which, for $p\geq 1$, reads:
$\mathbb{E}\left[|X+Y|^{p}\right]\leq 2^{p-1}\left(\mathbb{E}\left[|X|^{p}\right]+\mathbb{E}\left[|Y|^{p}\right]\right).$
That inequality with $X=\omega$ and $Y=-1$ delivers $q((\omega-1)^{p})\leq 2^{p-1}(1+q(\omega^{p}))$. This establishes (13).
For the non-centred moment, we proceed as follows:
$\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1+1|^{p}\right]=\sum_{k=0}^{p}{p\choose k}\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1|^{k}\right],$
then using Hölder’s inequality, this is less than
$\sum_{k=0}^{p}{p\choose k}\mathbb{E}\left[|\hat{Z}(\mathbf{x})-1|^{p}\right]^{k/p}\leq\sum_{k=0}^{p}{p\choose k}\left((p-1)^{p}N^{-p/2}2^{p-1}(1+q(\omega^{p}))\right)^{k/p}.$
From the binomial theorem, $\sum_{k=0}^{p}{p\choose k}a^{k}=(a+1)^{p}$, we obtain
$\mathbb{E}\left[|\hat{Z}(\mathbf{x})|^{p}\right]\leq\left(1+(p-1)N^{-1/2}2^{1-1/p}(1+q(\omega^{p}))^{1/p}\right)^{p}.$
This bound goes to one as $N\to\infty$. ∎
Proof of Lemma 1.1.
Using Markov’s inequality and Proposition 1.1, we have:
$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+z\right)$ $=\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})-1\geq z\right)$ (49)
$\leq\frac{\mathbb{E}_{\bar{q}}\left[\left|\hat{Z}(\mathbf{x})-1\right|^{p}\right]}{z^{p}}$ (50)
$\leq\frac{M(p)N^{-p/2}}{z^{p}}.$ (51)
The second line uses Markov’s inequality, and the third line applies Proposition 1.1. ∎
A.2 Proofs of Section 2
A.2.1 Proof of Theorem 2.1
We start with a technical result on the inverse moments of averages, which may be well-known.
Proposition A.1.
Let $r\geq 1$, $(x_{j})_{j\geq 0}$ a sequence of i.i.d. random variables with distribution $q$ on $\mathbb{X}$, and suppose that $\omega:\mathbb{X}\to(0,\infty)$ such that $q(\omega^{-\eta})<\infty$ for some $\eta>0$. Write $\omega_{j}=\omega(x_{j})$ for all $j=1,\ldots,N$. Then, for $N>\lfloor r/\eta\rfloor+1$, we have that
$\mathbb{E}\left[\left(\frac{N}{\omega_{1}+\cdots+\omega_{N}}\right)^{r}\right]\leq 2^{r}q(\omega^{-\eta})^{r/\eta}<\infty.$
Proof.
Let $\hat{W}=\frac{1}{N}(\omega_{1}+\cdots+\omega_{N})$. We will proceed by splitting the variables into blocks of size $j$ for $r/\eta\leq j\leq N$, which is possible by assumption, as follows: for $k\leq\lfloor N/j\rfloor$ we define
$\hat{W}^{j}_{k}:=\frac{1}{j}\left(\omega_{kj+1}+\cdots+\omega_{(k+1)j}\right)\qquad\text{and}\qquad\hat{W}^{j}_{\lfloor N/j\rfloor+1}:=\frac{1}{j}\left(\omega_{\lfloor N/j\rfloor j+1}+\cdots+\omega_{N}\right),$
where the final block may have fewer than $j$ elements. We lower bound $\hat{W}$ by dropping the last block if it has length strictly less than $j$,
$\hat{W}$ $\geq\frac{\hat{W}^{j}_{1}+\cdots+\hat{W}^{j}_{\left\lfloor\frac{N}{j}\right\rfloor}}{\frac{N}{j}}=\frac{\hat{W}^{j}_{1}+\cdots+\hat{W}^{j}_{\left\lfloor\frac{N}{j}\right\rfloor}}{\left\lfloor\frac{N}{j}\right\rfloor}\cdot\frac{\left\lfloor\frac{N}{j}\right\rfloor}{\frac{N}{j}}=:\frac{\left\lfloor\frac{N}{j}\right\rfloor}{\frac{N}{j}}\cdot\widetilde{W}.$
Since the mapping $m:x\mapsto 1/x^{r}$ is monotone decreasing and convex (we assumed $r\geq 1$), we have:
$\mathbb{E}\left[\hat{W}^{-r}\right]$ $=\mathbb{E}\left[m(\hat{W})\right]\leq\left(\frac{\frac{N}{j}}{\left\lfloor\frac{N}{j}\right\rfloor}\right)^{r}\mathbb{E}\left[m(\widetilde{W})\right]$
$\leq\left(1+\frac{1}{\left\lfloor\frac{N}{j}\right\rfloor}\right)^{r}\cdot\frac{1}{\left\lfloor\frac{N}{j}\right\rfloor}\sum_{k=1}^{\left\lfloor\frac{N}{j}\right\rfloor}\mathbb{E}[m(\hat{W}^{j}_{k})]$
$\leq\frac{2^{r}}{\left\lfloor\frac{N}{j}\right\rfloor}\sum_{k=1}^{\left\lfloor\frac{N}{j}\right\rfloor}\mathbb{E}\left[m\left(\hat{W}^{j}_{k}\right)\right].$
To proceed, we utilize the arithmetic-geometric mean inequality, which states that for non-negative numbers $a_{1},a_{2},\ldots,a_{j}$:
$\frac{a_{1}+a_{2}+\cdots+a_{j}}{j}\geq\left(a_{1}\cdot a_{2}\cdots a_{j}\right)^{\frac{1}{j}}.$
Applying this inequality, we obtain under the assumption that $q(\omega^{-\eta})<\infty$,
$\mathbb{E}\left[\left(\frac{j}{\omega_{1}+\dots+\omega_{j}}\right)^{r}\right]$ $\leq\mathbb{E}\left[\prod_{k=1}^{j}\left(\frac{1}{\omega_{k}}\right)^{r/j}\right]=\left(q(\omega^{-\frac{r}{j}})\right)^{j}$
$\leq q(\omega^{-\eta})^{j\frac{r}{j\eta}}=q(\omega^{-\eta})^{\frac{r}{\eta}}<\infty,$
where we have used Hölder’s inequality with the exponent $r^{\prime}=\eta j/r\geq 1$, by the choice of $j\geq r/\eta$. This yields the desired result. ∎
Proof of Theorem 2.1.
We first write the rescaled bias of normalized importance sampling as
$N\times\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[\hat{F}(\mathbf{x})-\pi(f)\right]$ $=\mathbb{E}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})(f(x_{n})-\pi(f))}{\sum_{n=1}^{N}\omega(x_{n})/N}\right]$ (52)
$=N\mathbb{E}\left[\frac{\omega(x_{1})(f(x_{1})-\pi(f))}{\sum_{n=1}^{N}\omega(x_{n})/N}\right]\quad\text{by identity in distribution}$ (53)
$=N\mathbb{E}\left[\frac{\omega(x_{1})(f(x_{1})-\pi(f))}{\sum_{n=2}^{N}\omega(x_{n})/N}\right]$ (54)
$+N\mathbb{E}\left[\omega(x_{1})(f(x_{1})-\pi(f))\left\{\frac{1}{\sum_{n=1}^{N}\omega(x_{n})/N}-\frac{1}{\sum_{n=2}^{N}\omega(x_{n})/N}\right\}\right].$ (55)
By independence and $\mathbb{E}[\omega(x_{1})f(x_{1})]=\pi(f)$, the first expectation is zero. For the second term,
$\frac{1}{\sum_{n=1}^{N}\omega(x_{n})/N}-\frac{1}{\sum_{n=2}^{N}\omega(x_{n})/N}=\frac{-\omega(x_{1})/N}{(\sum_{n=1}^{N}\omega(x_{n})/N)(\sum_{n=2}^{N}\omega(x_{n})/N)}.$ (56)
Thus, we can write
$N\times\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[\hat{F}(\mathbf{x})-\pi(f)\right]$ $=-N\mathbb{E}\left[\frac{\omega(x_{1})^{2}(f(x_{1})-\pi(f))/N}{(\sum_{n=1}^{N}\omega(x_{n})/N)(\sum_{n=2}^{N}\omega(x_{n})/N)}\right],$ (57)
and we further re-use (56) so that only $x_{j}$’s with $j\neq 1$ appear in the denominator of the leading term:
$N\times\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[\hat{F}(\mathbf{x})-\pi(f)\right]$ $=-\mathbb{E}\left[\frac{\omega(x_{1})^{2}(f(x_{1})-\pi(f))}{(\sum_{n=2}^{N}\omega(x_{n})/N)^{2}}\right]$ (58)
$-\mathbb{E}\left[\frac{-\omega(x_{1})}{(\sum_{n=1}^{N}\omega(x_{n})/N)(\sum_{n=2}^{N}\omega(x_{n})/N)}\times\frac{\omega(x_{1})^{2}(f(x_{1})-\pi(f))/N}{\sum_{n=2}^{N}\omega(x_{n})/N}\right].$ (59)
Having different $x_{j}$’s in the numerator and denominator, and using their independence, the leading term in (58) is $-q(\omega^{2}\cdot(f-\pi(f)))\mathbb{E}[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-2}]$.
Let $T_{N}=N^{-1}\sum_{n=2}^{N}\omega(x_{n})$. By the strong law of large numbers, $T_{N}^{-2}\xrightarrow{a.s.}1$ as $N\to\infty$. To strengthen this to convergence in $L^{1}$ of $T_{N}^{-2}$ to $1$, we use uniform integrability, e.g. Billingsley (1999), Theorem 3.5. A criterion for uniform integrability is (3.18) in Billingsley (1999), which is satisfied here since $\sup_{N}\mathbb{E}[T_{N}^{-3}]<\infty$ using $q(\omega^{-\eta})<\infty$ and Proposition A.1 with $r=3$, thus requiring $N>\lfloor 3/\eta\rfloor+1$.
It remains to show that the term in (59) goes to zero as $N\to\infty$. First we use the positivity of $\omega$ and the independence of $x_{j}$’s to get
$\mathbb{E}\left[\frac{\omega(x_{1})^{3}(f(x_{1})-\pi(f))/N}{(\sum_{n=1}^{N}\omega(x_{n})/N)(\sum_{n=2}^{N}\omega(x_{n})/N)^{2}}\right]$ $\leq\mathbb{E}\left[\left|\frac{\omega(x_{1})^{3}(f(x_{1})-\pi(f))/N}{(\sum_{n=2}^{N}\omega(x_{n})/N)^{3}}\right|\right]$ (60)
$=\frac{1}{N}\cdot\mathbb{E}\left[\left|\omega(x_{1})^{3}(f(x_{1})-\pi(f))\right|\right]\mathbb{E}\left[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-3}\right].$ (61)
The first expectation is finite by assumption. Using Proposition A.1, $\mathbb{E}\left[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-3}\right]\leq(N/(N-1))^{3}2^{3}q(\omega^{-\eta})^{3/\eta}$ when $N>\lfloor 3/\eta\rfloor+1$. Thus, this term in (59) behaves as a constant divided by $N$. ∎
A.2.2 Proof of Theorem 2.3
Proof of Theorem 2.3.
We write the IS estimator: $(\sum_{n=1}^{N}f(x_{n})\omega(x_{n}))/(\sum_{n=1}^{N}\omega(x_{n}))$, where $x_{1},\ldots,x_{n}$ are i.i.d. from $q$. We write the average weight: $q^{N}(\omega):=\sum_{n=1}^{N}\omega(x_{n})/N$.
First, it is enough to consider the case where the test function $f$ is non-negative. Indeed, for a general function $f$ we write $f=f_{+}-f_{-}$ where $f_{+}(x):=\max\{f(x),0\}$ and $f_{-}(x):=-\min\{f(x),0\}$. Then
$\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\pi(f)\right\rvert=\left\lvert\frac{q^{N}(f_{+}\omega)-q^{N}(f_{-}\omega)}{q^{N}(\omega)}-(\pi(f_{+})-\pi(f_{-}))\right\rvert\leq\left\lvert\frac{q^{N}(f_{+}\omega)}{q^{N}(\omega)}-\pi(f_{+})\right\rvert+\left\lvert\frac{q^{N}(f_{-}\omega)}{q^{N}(\omega)}-\pi(f_{-})\right\rvert.$
Using $(a+b)^{s}\leq 2^{s-1}(a^{s}+b^{s})$, and applying the result for non-negative functions $f_{+}$ and $f_{-}$ separately, we obtain the result for general $f$. Thus, we now assume that $f$ takes non-negative values.
We write the absolute error between the IS estimator with the target $\pi(f)=q(f\omega)$ in two different ways. The first is:
$\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert\leq\max_{1\leq i\leq N}f(x_{i})+q(f\omega).$ (62)
The second is:
$\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert\leq\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\frac{q(f\omega)}{q^{N}(\omega)}\right\rvert+q(f\omega)\left\lvert\frac{1}{q^{N}(\omega)}-1\right\rvert.$
Now we consider two cases: 1) $\lvert q^{N}(\omega)-1\rvert>0.5$, 2) $\lvert q^{N}(\omega)-1\rvert\leq 0.5$. We will separately bound the expected error under the two cases using the two inequalities above.
We start with the first case, and we assume $r<\infty$. First, we use $(a+b)^{s}\leq 2^{s-1}(a^{s}+b^{s})$ to write
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$ $\leq\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})+q(f\omega)\right)^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$
$\leq 2^{s-1}\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})\right)^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]\break$
$+2^{s-1}\left(q(f\omega)\right)^{s}\mathbb{P}[\lvert q^{N}(\omega)-1\rvert>0.5].$
The second term leads to a bound in $N^{-s/2}$ using Markov’s inequality as in Lemma 1.1, since $q(\omega^{s})<\infty$ under the assumptions. The first term is dealt with first using Hölder’s inequality with exponents $r/s$ and $(1-s/r)^{-1}$,
$\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})\right)^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$
$\leq\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})\right)^{r}\right]^{s/r}\times\mathbb{P}[\lvert q^{N}(\omega)-1\rvert>0.5]^{1-s/r}$
$\leq q(f^{r})^{s/r}N^{s/r}\cdot C\cdot N^{-0.5p(1-s/r)},$
for a constant $C$. The last inequality uses the fact that $\mathbb{E}[(\max_{1\leq i\leq N}f(x_{i}))^{r}]\leq N\mathbb{E}[f(x_{1})^{r}]$, and Markov’s inequality using $q(\omega^{p})<\infty$. Given $s\leq pr/(p+r+2)$, the exponent of $N$ satisfies
$\frac{s}{r}-\frac{p(r-s)}{2r}=\frac{2s+ps-pr}{2r}\leq\frac{-s}{2},$
using $s\leq pr/(p+r+2)\Leftrightarrow-pr\leq-s(p+r+2)$. Altogether we arrive at
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]\leq CN^{-s/2},$
for another constant $C$.
In the case $r=\infty$, we can directly write
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-q(f\omega)\right\rvert^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$ $\leq\mathbb{E}\left[\left(\max_{1\leq i\leq N}f(x_{i})+q(f\omega)\right)^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert>0.5)\right]$
$\leq 2^{s}\lvert f\rvert_{\infty}^{s}\mathbb{P}[\lvert q^{N}(\omega)-1\rvert>0.5]$
$\leq 2^{s}C\lvert f\rvert_{\infty}^{s}N^{-0.5p}\leq CN^{-0.5s},$
using $s\leq\min\{p,r\}\leq p$ in the last line, and changing the value of $C$ between inequalities.
For the case $|q^{N}(\omega)-1|\leq 0.5$,
$\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\pi(f)\right\rvert\mathds{1}(\lvert q^{N}(\omega)-1\rvert\leq 0.5)$ $\leq 2\lvert q^{N}(f\omega)-\pi(f)\rvert+\pi(f)\left\lvert\frac{q^{N}(\omega)-1}{q^{N}(\omega)}\right\rvert$
$\leq 2\lvert q^{N}(f\omega)-\pi(f)\rvert+2\pi(f)\left\lvert q^{N}(\omega)-1\right\rvert.$
Therefore
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\pi(f)\right\rvert^{s}\mathds{1}(\lvert q^{N}(\omega)-1\rvert<0.5)\right]$ $\leq C\left(\mathbb{E}[\lvert q^{N}(f\omega)-\pi(f)\rvert^{s}]+\mathbb{E}[\left\lvert q^{N}(\omega)-1\right\rvert^{s}]\right)$
$\leq CN^{-s/2},$
for some constant $C$ that changes at each line. The first term is $O(N^{-s/2})$ with a reasoning similar to that in the proof of Proposition 1.1, since $q^{N}(f\omega)$ is the sum of $N$ i.i.d. random variables with mean $q(f\omega)$ and $s$ finite moments, since $s\leq pr/(p+r+2)\leq pr/p+r$. Putting everything together gives
$\mathbb{E}\left[\left\lvert\frac{q^{N}(f\omega)}{q^{N}(\omega)}-\pi(f)\right\rvert^{s}\right]\leq CN^{-s/2}.$
A.3 Proofs of Section 3
We prove Theorem 3.1. We assume that both target and proposal distributions admit densities with respect to a measure $\lambda$. Although we will express all subsequent notations using integration, this should be interpreted as summation when the space is discrete and $\lambda$ represents the counting measure. The rejection probability at $\mathbf{x}$ is denoted by
$r(\mathbf{x})=\int_{\mathbf{z}\neq\mathbf{x}}\left(1-\min\left(1,\frac{\hat{Z}(\mathbf{z})}{\hat{Z}(\mathbf{x})}\right)\bar{q}(\mathbf{z})\right)\lambda(\mathrm{d}\mathbf{z}).$
That definition only considers the probability of moves to states different than $\mathbf{x}$ that are rejected. We will use the following fact: at every iteration, for each chain one of the following three events occurs: 1) a proposal to a different state is accepted, 2) a proposal to a different state is rejected, 3) a proposal is made to the current state (and systematically accepted). In a continuous state space with an atomless measure $\lambda$, the last event occurs with probability zero.
We first prove a lemma that describes the coupling time $\tau$.
Lemma A.1.
Assuming $\omega(\mathbf{x})\geq\omega(\mathbf{y})$, we have the following facts:
  • Let $\tau_{0}$ be the first time when the $\mathbf{x}$-chain moves to a different state. Then $\tau\leq\tau_{0}$, i.e. the chains meet at $\tau_{0}$ or earlier.
  • Let $\tau_{1}$ be the first time when a common proposal is $\mathbf{x}$. Then $\tau\leq\tau_{1}$, i.e. the chains meet at $\tau_{1}$ or earlier.
  • The meeting time satisfies $\tau=\min\{\tau_{0},\tau_{1}\}$.
Proof of Lemma A.1.
The first two observations can be proven by induction, once we recognize that the common draws coupling of Algorithm 4 implies $\omega(\mathbf{x}_{t})\geq\omega(\mathbf{y}_{t})$ for all $t\geq 0$. Regarding the last observation, for every $t<\min\{\tau_{0},\tau_{1}\}$, the $\mathbf{x}$-chain must have rejected moves to a different state than $\mathbf{x}$ at each iteration up to $t$. In that situation, the $\mathbf{x}$-chain is still at $\mathbf{x}$, while the $\mathbf{y}$-chain never proposed a move to $\mathbf{x}$ and thus $\mathbf{x}_{t}=\mathbf{x}\neq\mathbf{y}_{t}$, as claimed. ∎
Now we calculate the tail probability of $\tau$.
Lemma A.2.
For all $t\geq 1$, $|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\leq\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)=\max(r(\mathbf{x}),r(\mathbf{y}))^{t}.$
Proof of Lemma A.2.
The inequality in the statement is the celebrated coupling inequality. For the equality, we assume $\omega(\mathbf{x})\geq\omega(\mathbf{y})$ without loss of generality, which implies $r(\mathbf{x})\geq r(\mathbf{y})$. By Lemma A.1, the event $\{\tau>t\}$ is equivalent to $\{\min\{\tau_{0},\tau_{1}\}>t\}$. The latter event corresponds to the event: “the $\mathbf{x}$-chain proposes to move to a different state but gets rejected at each of the first $t$ iterations”. Then its probability is $r(\mathbf{x})^{t}$, since $r(\mathbf{x})$ is the probability of a failed attempt to move to a different state. ∎
It remains to show the following lower bound.
Lemma A.3.
For all $t\geq 1$, $|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\geq\mathbb{P}_{\mathbf{x},\mathbf{y}}(\tau>t)$.
Proof of Lemma A.3.
Again, we assume $\omega(\mathbf{x})\geq\omega(\mathbf{y})$ without loss of generality. The definition of total variation distance as a supremum over measurable sets implies $|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)|_{\text{TV}}\geq P^{t}(\mathbf{x},\{\mathbf{x}\})-P^{t}(\mathbf{y},\{\bf{x}\})$, considering the set $\{\mathbf{x}\}$.
Under the distribution of the coupled chains, we can write $P^{t}(\mathbf{x},\{\mathbf{x}\})-P^{t}(\mathbf{y},\{\bf{x}\})$ as $\mathbb{P}(\mathbf{x}_{t}=\mathbf{x})-\mathbb{P}(\mathbf{y}_{t}=\mathbf{x})$. Now we decompose each probability according to $\tau$ being greater or less than $t$, for any $t\geq 1$:
$\mathbb{P}\left(\mathbf{x}_{t}=\bf x\right)-\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x}\right)=\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x};\tau>t\right)+\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x};\tau\leq t\right)-\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x};\tau>t\right)-\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x};\tau\leq t\right).$
We simplify with the following observations.
  • Under the event $\tau>t$: we have $\mathbf{x}_{t}=\mathbf{x}$; otherwise, the $\mathbf{x}$-chain would have successfully moved to a new state jointly with the $\mathbf{y}$-chain implying $\tau\leq t$ by Lemma A.1. Therefore,
    $\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x};\tau>t\right)=\mathbb{P}\left(\tau>t\right)\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x}\mid\tau>t\right)=\mathbb{P}\left(\tau>t\right).$
    Meanwhile, under that event we have $\mathbf{y}_{t}\neq\mathbf{x}$; otherwise, the $\mathbf{y}$-chain must have proposed a move to $\mathbf{x}$ at or before time $t$, and that would have resulted in a meeting by Lemma A.1. Therefore,
    $\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x};\tau>t\right)=0.$
  • Under the event $\tau\leq t$: we have $\mathbf{x}_{t}=\mathbf{y}_{t}$, therefore $\mathbb{P}(\mathbf{x}_{t}=\mathbf{x};\tau\leq t)=\mathbb{P}(\mathbf{y}_{t}=\mathbf{x};\tau\leq t)$.
Putting these together, we conclude that $\mathbb{P}\left(\mathbf{x}_{t}=\mathbf{x}\right)-\mathbb{P}\left(\mathbf{y}_{t}=\mathbf{x}\right)=\mathbb{P}\left(\tau>t\right)$. ∎
Theorem 3.1 is obtained by combining Lemmas A.2 and A.3.
A.4 Proofs of Section 4
A.4.1 Proof of Proposition 4.1
Let $t\geq 1$. The event $\{\tau>t\}$ only occurs when Algorithm 5 enters its while loop, in which case we must have that 1) $\mathbf{x}_{1}=\mathbf{x}$, 2) $\hat{Z}(\mathbf{x})>\hat{Z}(\mathbf{y}_{0})$, and 3) the first generated Uniform variable was greater than $\hat{Z}(\mathbf{y}_{0})/\hat{Z}(\mathbf{x})$. Thus,
$\mathbb{P}(\tau>t)=\iint\mathbb{P}_{\mathbf{x},\mathbf{y}_{0}}\left(\tau>t\right)\left(1-\min\left\{1,\frac{\hat{Z}(\mathbf{y}_{0})}{\hat{Z}(\mathbf{x})}\right\}\right)\mathds{1}\left(\hat{Z}(\mathbf{y}_{0})<\hat{Z}(\mathbf{x})\right)\mathds{1}(\mathbf{y}_{0}\neq\mathbf{x})\bar{q}(\mathrm{d}\mathbf{x})\bar{q}(\mathrm{d}\mathbf{y}_{0}).$ (63)
The quantity $\mathbb{P}_{\mathbf{x},\mathbf{y}_{0}}(\tau>t)$ in the event $\hat{Z}(\mathbf{y}_{0})<\hat{Z}(\mathbf{x})$ is equal to $r(\mathbf{x})^{t}$, as in Theorem 3.1. By upper-bounding the other terms by one and integrating with respect to $\bar{q}(\mathrm{d}\mathbf{y}_{0})$, we obtain the upper bound
$\mathbb{P}(\tau>t)\leq\int(r(\mathbf{x}))^{t}\bar{q}(\mathrm{d}\mathbf{x})=\mathbb{E}_{\bar{q}}\left[(r(\mathbf{x}))^{t}\right].$ (64)
A.4.2 Proof of Proposition 4.2
We prove Proposition 4.2 by first splitting the expectation according to whether $\hat{Z}(\mathbf{x})$ is less than or greater than 2:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\right]=\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\leq 2)\right]+\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})>2)\right].$ (65)
We then proceed through a series of lemmas to bound each term. The following lemmas are used to handle the case when $\hat{Z}(\mathbf{x})>2$:
Lemma A.4.
Under Assumption 1 and $q(\omega^{p})<\infty$ for any $p>1$, the rejection probability (19) is upper bounded as follows, for any $\theta\in[0,1]$:
$r(\mathbf{x})\leq 1-\min\left\{1,\frac{\theta}{\hat{Z}(\mathbf{x})}\right\}c_{p}(\theta),\quad\text{with}\quad c_{p}(\theta)=\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}\in[0,1].$ (66)
Proof.
Let $\theta\in[0,1]$. We start with an $L^{p}$-version of Paley-Zygmund inequality, as in page 2705, equation (12) of Petrov (2007) with $r=1$. If $W$ is a non-negative random variable and $p>1$, then
$\mathbb{P}\left(W>\theta\mathbb{E}[W]\right)\geq\frac{(1-\theta)^{p/(p-1)}\left(\mathbb{E}[W]\right)^{p/(p-1)}}{\left(\mathbb{E}[W^{p}]\right)^{1/(p-1)}}.$ (67)
Indeed, for any $b>0$, Hölder’s inequality implies
$\mathbb{E}[W]$ $=\mathbb{E}[W\mathds{1}(W>b)]+\mathbb{E}[W\mathds{1}(W\leq b)]$
$\leq\mathbb{P}(W>b)^{(1-1/p)}\mathbb{E}[W^{p}]^{1/p}+b.$
Re-arranging with $b=\theta\mathbb{E}[W]$ implies (67). We apply this to $\hat{Z}(\mathbf{x})$, under Assumption 1:
$\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})>\theta\right)\geq\frac{(1-\theta)^{p/(p-1)}}{\left(\mathbb{E}_{\bar{q}}\left[\left(\hat{Z}(\mathbf{x})\right)^{p}\right]\right)^{1/(p-1)}}\geq\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}.$ (68)
The latter inequality comes from Jensen’s, since $z\mapsto z^{p}$ is convex since $p>1$:
$\mathbb{E}_{\bar{q}}\left[\left(\hat{Z}(\mathbf{x})\right)^{p}\right]\leq\mathbb{E}_{\bar{q}}\left[\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n})^{p}\right]=q(\omega^{p}).$ (69)
Inequality (68) implies that
$\displaystyle\int\min\left\{1,\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})}\right\}\bar{q}(\mathrm{d}\mathbf{x}^{\star})$ =$=\int_{\{\mathbf{x}^{\star}:\hat{Z}(\mathbf{x}^{\star})\leq\theta\}}\min\left\{1,\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})}\right\}\bar{q}(\mathrm{d}\mathbf{x}^{\star})$
+$\quad+\int_{\{\mathbf{x}^{\star}:\hat{Z}(\mathbf{x}^{\star})>\theta\}}\min\left\{1,\frac{\hat{Z}(\mathbf{x}^{\star})}{\hat{Z}(\mathbf{x})}\right\}\bar{q}(\mathrm{d}\mathbf{x}^{\star})$
$\geq 0+\min\left\{1,\frac{\theta}{\hat{Z}(\mathbf{x})}\right\}\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x}^{\star})>\theta\right)$
$\geq\min\left\{1,\frac{\theta}{\hat{Z}(\mathbf{x})}\right\}\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}.$
This yields the desired result. ∎
Lemma A.5.
Under Assumptions 1-2, there exists a constant $C>0$ such that for all $t\geq 1$, $N\geq 1$,
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})>2)\right]\leq\frac{C}{N^{p/2}t^{p}}.$ (70)
Proof.
We split the expectation into two parts:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})>2\right)\right]$ $=\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\in(2,1+t)\right)\right]+\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\geq 1+t\right)\right].$ (71)
For $\{\hat{Z}(\mathbf{x})\geq 1+t\}$, we directly apply Lemma 1.1:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\geq 1+t)\right]$ $\leq\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+t\right)$ (72)
$\leq\frac{M(p)}{N^{p/2}t^{p}}.$ (73)
For $\{\hat{Z}(\mathbf{x})\in(2,1+t)\}$, we use Lemma A.4 with $\theta=1/2$:
$r(\mathbf{x})$ $\leq 1-\frac{c_{p}(1/2)}{2\hat{Z}(\mathbf{x})}\leq\exp\left(-\frac{c_{p}(1/2)}{2\hat{Z}(\mathbf{x})}\right).$ (74)
Let $c=c_{p}(1/2)/4$. Then using the fact that $\hat{Z}(\mathbf{x})>2$ implies that $\hat{Z}(\mathbf{x})\leq 2(\hat{Z}(\mathbf{x})-1)$, we have:
$r(\mathbf{x})^{t}$ $\leq\exp\left(-\frac{2ct}{\hat{Z}(\mathbf{x})}\right)\leq\exp\left(-\frac{ct}{\hat{Z}(\mathbf{x})-1}\right).$ (75)
We introduce the sets $A_{k}=\left[t/(k+1),t/k\right]$ for $k\geq 1$, so that $\cup_{k=1}^{\infty}A_{k}=[0,t]$ which contains $[1,t]$. Using the result of Lemma 1.1, we obtain the bound:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}(\hat{Z}(\mathbf{x})\in(2,1+t))\right]$ $\leq\mathbb{E}_{\bar{q}}\left[\exp\left(-\frac{ct}{\hat{Z}(\mathbf{x})-1}\right)\mathds{1}\left(\hat{Z}(\mathbf{x})-1\in(1,t)\right)\right]$ (76)
$\leq\sum_{k=1}^{\infty}\mathbb{E}_{\bar{q}}\left[\exp\left(-\frac{ct}{\hat{Z}(\mathbf{x})-1}\right)\mathds{1}\left(\hat{Z}(\mathbf{x})-1\in A_{k}\right)\right]$ (77)
$\leq\sum_{k=1}^{\infty}\exp(-ck)\mathbb{P}_{\bar{q}}\left(\hat{Z}(\mathbf{x})\geq 1+t/(k+1)\right)$ (78)
$\leq\sum_{k=1}^{\infty}\exp(-ck)\frac{M(p)}{N^{p/2}}\left(\frac{k+1}{t}\right)^{p}.$ (79)
Let $S_{p}=\sum_{k=1}^{\infty}\exp(-ck)(k+1)^{p}$, which is finite. Then:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\in(2,1+t)\right)\right]$ $\leq\frac{M(p)S_{p}}{N^{p/2}t^{p}}.$ (80)
Combining the bounds for both parts, we get:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})>2\right)\right]$ $\leq\frac{M(p)}{N^{p/2}t^{p}}+\frac{M(p)S_{p}}{N^{p/2}t^{p}}$ (81)
$\leq\frac{M(p)(1+S_{p})}{N^{p/2}t^{p}}.$ (82)
Setting the new constant $C:=M(p)\left(1+S_{p}\right)$ completes the proof. ∎
Now, we turn our attention to controlling the expectation when $\hat{Z}(\mathbf{x})\leq 2$.
Lemma A.6.
Fix $p\geq 2$ and let $\beta_{p}$ be defined as in (24). There exist constants $A_{p},B_{p}>0$, depending only on $p$ and $q(\omega^{p})$, such that for all $N\geq 1$, for all $t\geq 1$, the following holds:
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]\leq\left[\frac{A_{p}}{N^{\frac{t\wedge p}{2}}}+\frac{B_{p}}{N^{p/2}}\right]\beta_{p}^{t}.$ (83)
Proof.
We abuse notation to write $r$ as a function of the value $z$ taken by $\hat{Z}(\mathbf{x})$, instead of a function of $\mathbf{x}$, in various places in this proof. First notice that $r(z)$ is increasing in $z$. We thus have that for $t\geq 1$ that
$\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]$
$\leq r(2)^{t-t\wedge p}\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t\wedge p}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right].$
We first consider the second factor. We have for any $\alpha\in(0,1)$,
$\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t\wedge p}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]$
$\leq\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t\wedge p}\mathds{1}\left(1-\alpha\leq\hat{Z}(\mathbf{x})\leq 2\right)\right]+r(2)^{t\wedge p}\bar{q}\left\{|\hat{Z}(\mathbf{x})-1|\geq\alpha\right\}.$ (84)
That is because $\{\hat{Z}(\mathbf{x})\leq 1-\alpha\}\subset\{|\hat{Z}(\mathbf{x})-1|\geq\alpha\}$, and $r(z)\leq r(2)$ for $z\leq 2$.
At this point, notice that by Lemma A.4 with $\theta=1/2$ we have that $r(2)\leq 1-c_{p}(1/2)/4=\beta_{p}$, where $\beta_{p}$ is defined in (24). Also notice that
$r(z)$ $=1-\int\min\left\{1,\frac{z^{\ast}}{z}\right\}\bar{q}(dz^{\ast})=\int_{z^{\ast}=0}^{\infty}\bar{q}(dz^{\ast})-\int_{z^{\ast}=0}^{z}\frac{z^{\ast}}{z}\bar{q}(dz^{\ast})-\int_{z^{\ast}=z}^{\infty}\bar{q}(dz^{\ast})$
$=\int_{z^{\ast}=0}^{z}\bar{q}(dz^{\ast})-\int_{z^{\ast}=0}^{z}\frac{z^{\ast}}{z}\bar{q}(dz^{\ast})=\int_{z^{\ast}=0}^{z}\left(\frac{z-z^{\ast}}{z}\right)\bar{q}(dz^{\ast}).$
Returning to our calculation regarding the first term in (84),
$\mathbb{E}_{\bar{q}}\left[r(\hat{Z}(\mathbf{x}))^{t\wedge p}\mathds{1}\left(1-\alpha\leq\hat{Z}(\mathbf{x})\leq 2\right)\right]$
$=\mathbb{E}_{\bar{q}}\left[\left(\frac{1}{\hat{Z}(\mathbf{x})}\int_{z^{\ast}=0}^{\hat{Z}(\mathbf{x})}(\hat{Z}(\mathbf{x})-z^{\ast})\bar{q}(dz^{\ast})\right)^{t\wedge p}\mathds{1}\left(1-\alpha\leq\hat{Z}(\mathbf{x})\leq 2\right)\right]$
$\leq\frac{\bar{q}\{[0,\hat{Z}(\mathbf{x})]\}^{t\wedge p}}{(1-\alpha)^{t\wedge p}}\mathbb{E}_{\bar{q}}\left[\left(\int_{z^{\ast}=0}^{\hat{Z}(\mathbf{x})}(\hat{Z}(\mathbf{x})-z^{\ast})\frac{\bar{q}(dz^{\ast})}{\bar{q}\{[0,\hat{Z}(\mathbf{x})]\}}\right)^{t\wedge p}\mathds{1}\left(1-\alpha\leq\hat{Z}(\mathbf{x})\leq 2\right)\right]$
[…]
$\leq\frac{1}{(1-\alpha)^{t\wedge p}}\mathbb{E}_{\bar{q}}\left[\int_{z^{\ast}=0}^{\hat{Z}(\mathbf{x})}|\hat{Z}(\mathbf{x})-z^{\ast}|^{t\wedge p}\bar{q}(dz^{\ast})\right]$
$\leq\frac{1}{(1-\alpha)^{t\wedge p}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{\prime})\sim\bar{q}\otimes\bar{q}}\left[|\hat{Z}(\mathbf{x})-\hat{Z}(\mathbf{x}^{\prime})|^{t\wedge p}\right]\leq\frac{1}{(1-\alpha)^{t\wedge p}}\frac{A_{p}\bar{q}(\omega^{p})}{N^{\frac{t\wedge p}{2}}},$
for a constant $A_{p}$ depending only on $p$. The first inequality comes from $\hat{Z}(\mathbf{x})^{-1}\leq(1-\alpha)^{-1}$ on the event of interest, the second inequality is from Jensen’s since the function $u\mapsto u^{t\wedge p}$ is convex, the third inequality is from $\bar{q}(A)\leq 1$ and the indicator being smaller than one, the fourth is obtained by completing the integral over all $z^{\ast}\in(0,\infty)$, and the last is from a reasoning similar to the proof of Proposition 1.1, or by direct application of Minkowski’s inequality and Proposition 1.1.
Overall, choosing $\alpha=1/2$ we have that
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t\wedge p}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]$ $\leq 2^{t\wedge p}A_{p}\bar{q}(\omega^{p})N^{-t\wedge p/2}+r(2)^{t\wedge p}\bar{q}\left\{|\hat{Z}(\mathbf{x})-1|\geq\alpha\right\}$
$\leq 2^{t\wedge p}A_{p}\bar{q}(\omega^{p})N^{-t\wedge p/2}+\beta_{p}^{t\wedge p}C_{p}N^{-p/2}2^{p},$
using Markov’s inequality as in Lemma 1.1. Finally, multiplying by $r(2)^{t-t\wedge p}$ we obtain
$\mathbb{E}_{\bar{q}}\left[r(\mathbf{x})^{t}\mathds{1}\left(\hat{Z}(\mathbf{x})\leq 2\right)\right]\leq\frac{\beta_{p}^{t}\beta_{p}^{-t\wedge p}2^{t\wedge p}A_{p}\bar{q}(\omega^{p})}{N^{\frac{t\wedge p}{2}}}+\frac{2^{p}C_{p}\beta_{p}^{t}}{N^{p/2}},$
and we note that, since $\beta_{p}\leq 1$, we have $\beta_{p}^{-t\wedge p}2^{t\wedge p}\leq\beta_{p}^{-p}2^{p}$, and thus we can define $A_{p}$ and $B_{p}$ to obtain Lemma A.6. ∎
Proof of Proposition 4.2.
We combine the bounds from Lemmas A.5 and A.6, and note that the two terms in the bound of Lemma A.6 can be bounded by $A_{p}\beta_{p}^{t}N^{-(t\wedge p)/2}$ for some constant $A_{p}$, which is not the same $A_{p}$ as in the statement of Lemma A.6. ∎
A.4.3 Proofs of Theorem 4.1 and Corollary 4.1
Proof of Theorem 4.1.
Under Assumption 1, the PIMH chain is $\bar{\pi}$-irreducible, and by construction it is aperiodic and $\bar{\pi}$-invariant, therefore $|\bar{q}P^{t}-\bar{\pi}|_{\mathrm{TV}}\to 0$ as $t\to\infty$ (Theorem 4 in Roberts & Rosenthal 2004). Thus, for any $t\geq 0$, by the triangle inequality,
$|\bar{q}P^{t}-\bar{\pi}|_{\mathrm{TV}}\leq\sum_{j=1}^{\infty}|\bar{q}P^{t+j}-\bar{q}P^{t+j-1}|_{\mathrm{TV}}.$ (85)
By the coupling representation of the TV distance, for any $t\geq 0$, $j\geq 1$,
$|\bar{q}P^{t+j}-\bar{q}P^{t+j-1}|_{\mathrm{TV}}\leq\mathbb{E}[\mathds{1}(\mathbf{x}_{t+j}\neq\mathbf{y}_{t+j-1})]=\mathbb{P}(\tau>t+j),$ (86)
where $(\mathbf{x}_{t})$ and $(\mathbf{y}_{t})$ are jointly generated by Algorithm 5. Under Assumption 2, Proposition 4.3 applies and thus the series $\sum_{j=1}^{\infty}\mathbb{P}(\tau>t+j)$ converges. Thus, by the dominated convergence theorem we may swap expectation and limit to write
$|\bar{q}P^{t}-\bar{\pi}|_{\mathrm{TV}}\leq\mathbb{E}\left[\sum_{j=1}^{\infty}\mathds{1}(\mathbf{x}_{t+j}\neq\mathbf{y}_{t+j-1})\right]=\mathbb{E}\left[\max(0,\tau-t-1)\right],$ (87)
for all $t\geq 0$. This is (27).
We may express the expectation of a non-negative variable as a series of survival probabilities:
$\mathbb{E}\left[\max\left(0,\tau-t-1\right)\right]=\sum_{s=1}^{\infty}\mathbb{P}\left(\max\left(0,\tau-t-1\right)\geq s\right).$
For any $t\geq 0,s\geq 1$, $\max(0,\tau-1-t)\geq s$ if and only if $\tau>s+t$. Under Assumption 2, Proposition 4.3 obtains
$\mathbb{P}\left(\tau>s+t\right)\leq CN^{-1/2}(s+t)^{-p}.$
The series $\sum_{s=1}^{\infty}(s+t)^{-p}$ can be bounded as follows:
$\sum_{s=1}^{\infty}(s+t)^{-p}$ $=\sum_{s=1+t}^{\infty}s^{-p}=(1+t)^{-p}+\sum_{s=t+2}^{\infty}s^{-p}$ (88)
[…]
$={(1+t)}^{-p}+\left[-\frac{x^{-p+1}}{p-1}\right]_{1+t}^{\infty}$ (90)
$={(1+t)}^{-p}+\frac{(1+t)^{-p+1}}{p-1}$ (91)
$=(1+t)^{-p+1}\left(\frac{1}{1+t}+\frac{1}{p-1}\right)$ (92)
$=(1+t)^{-p+1}\left(\frac{(1+t/p)p}{(1+t)(p-1)}\right)$ (93)
$\leq\frac{p}{(p-1)(1+t)^{p-1}},$ (94)
[…]
$|\bar{q}P^{t}-\bar{\pi}|_{\text{TV}}\leq\frac{Cp}{\sqrt{N}(p-1)(1+t)^{p-1}},$
which completes the proof. ∎
Proof of Corollary 4.1.
The proof starts with multiple applications of the triangle inequality, Theorem 3.1, $\max(a,b)\leq a+b$ for $a,b\geq 0$:
$\left|P^{t}(\mathbf{x},\cdot)-\bar{\pi}\right|_{\mathrm{TV}}$ $\leq\left|P^{t}(\mathbf{x},\cdot)-\bar{q}P^{t}\right|_{\mathrm{TV}}+\left|\bar{q}P^{t}-\bar{\pi}\right|_{\mathrm{TV}}$
$\leq\displaystyle\int\left|P^{t}(\mathbf{x},\cdot)-P^{t}(\mathbf{y},\cdot)\right|\bar{q}(\mathrm{d}\mathbf{y})+\left|\bar{q}P^{t}-\bar{\pi}\right|_{\mathrm{TV}}$
$=\displaystyle\int\max\left(r(\mathbf{x}),r(\mathbf{y})\right)^{t}\bar{q}(\mathrm{d}\mathbf{y})+\left|\bar{q}P^{t}-\bar{\pi}\right|_{\mathrm{TV}}$
$\leq\left(r(\mathbf{x})\right)^{t}+\mathbb{E}_{\bar{q}}[\left(r(\mathbf{y})\right)^{t}]+\left|\bar{q}P^{t}-\bar{\pi}\right|_{\mathrm{TV}}.$
The result then follows from Proposition 4.2 and Theorem 4.1. ∎
Proof of Proposition 4.4.
Similarly to the proof of Corollary 4.1, we start from
$|P^{t}(x,\cdot)-\pi|_{\mathrm{TV}}=|P^{t}(x,\cdot)-\pi P^{t}|_{\mathrm{TV}}\leq r(x)^{t}+\mathbb{E}_{y\sim\pi}[r(y)^{t}].$ (95)
Let $p>1$. We can apply Lemma A.4 to obtain
$r(y)\leq 1-\min\left\{1,\frac{\theta}{\omega(y)}\right\}\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}},$ (96)
and we set $\theta=1/2$, and $c=\frac{(1-\theta)^{p/(p-1)}}{q(\omega^{p})^{1/(p-1)}}$. Note that $c\leq 1$ as $q(\omega^{p})\geq q(\omega)^{p}=1$. We next bound the expected rejection probability as follows
$\mathbb{E}_{y\sim\pi}\left[r(y)^{t}\right]$ $\leq\mathbb{E}_{y\sim\pi}\left[\left(1-\min\left\{\frac{0.5}{\omega(y)},1\right\}c\right)^{t}\right]$ (97)
$\leq\mathbb{E}_{y\sim\pi}[(1-0.5c)^{t}I(\omega(y)\leq 1)]+\mathbb{E}_{y\sim\pi}\left[\left(1-\frac{0.5}{\omega(y)}c\right)^{t}I(\omega(y)\in[1,t])\right]+\mathbb{P}[\omega(y)\geq t]$ (98)
$\leq(1-0.5c)^{t}+\mathbb{E}_{y\sim\pi}\left[\left(1-\frac{0.5}{\omega(y)}c\right)^{t}I(\omega(y)\in[1,t])\right]+\frac{\tilde{C}}{t^{p-1}}$ (99)
$\leq(1-0.5c)^{t}+\mathbb{E}_{y\sim\pi}\left[\exp\{-Ct/\omega(y)\}I(\omega(y)\in[1,t])\right]+\frac{\tilde{C}}{t^{p-1}}.$ (100)
The second inequality follows by splitting the weight into $\omega\leq 1,\omega\in[1,t]$ and $\omega>t$. The third inequality employs Markov’s inequality and the assumption that $p>1$. The last inequality uses $\log(1+x)\leq x$ with $x=-0.5c/\omega(y)$, $C=0.5c$. Consider the three terms on the last line. The first term decays exponentially fast with $t$, the third term decays at the rate of $t^{-(p-1)}$. It remains to bound the second term.
Define $A_{k}:=[t/(k+1),t/k]$, then clearly $\cup_{k=1}^{\infty}A_{k}=[0,t]$. We bound the second term as follows:
$\mathbb{E}_{y\sim\pi}\left[\exp\{-Ct/\omega(y)\}I(\omega(y)\in[1,t])\right]$ $\leq\mathbb{E}_{y\sim\pi}\left[\exp\{-Ct/\omega(y)\}I(\omega(y)\in[0,t])\right]$
$=\sum_{k=1}^{\infty}\mathbb{E}_{y\sim\pi}\left[\exp\{-Ct/\omega(y)\}I(\omega(y)\in A_{k})\right]$
$\leq\sum_{k=1}^{\infty}\exp\{-Ct/(t/k)\}\mathbb{P}[\omega(y)\geq t/(k+1)]$
$\leq\sum_{k=1}^{\infty}\exp\{-Ck\}\frac{C^{\prime}(k+1)^{p-1}}{t^{p-1}}$
$=\frac{C^{\prime\prime}}{t^{p-1}}\sum_{k=1}^{\infty}\exp\{-Ck\}(k+1)^{p-1}$
$\leq\frac{C^{\prime\prime\prime}}{t^{p-1}}.$
The last inequality holds as $\sum_{k=1}^{\infty}\exp\{-Ck\}(k+1)^{p-1}<\infty$ (the terms inside the summation decay exponentially fast). This concludes the proof. ∎
A.4.4 Proof of the result in Example 3
To complement the upper bound in Corollary 4.1, we present an example where $q(\omega^{p})<\infty$, and $\left|P^{t}(\mathbf{x}_{0},\cdot)-\bar{\pi}\right|_{\text{TV}}=\tilde{\Omega}(t^{-(p-1)})$ for some $\mathbf{x}_{0}$. Here $\Omega$ hides constants that may depend on $p$, and $\tilde{\Omega}$ indicates that we are disregarding polylogarithmic factors with respect to $t$. We set $N=1$ here as the focus is on the rate in $t$, and we revert to IMH notation for simplicity.
Let us consider $\pi(x):=Z_{\pi}x^{-p}$ on $[2,\infty)$, and $q(x):=Z_{q}\log^{2}(x)x^{-(p+1)}$ on $[2,\infty)$. In this case $\omega(x)=(Z_{\pi}/Z_{q})(x/\log^{2}(x))$. We can check:
$q(\omega^{p})=\pi(\omega^{p-1})=(Z_{\pi}/Z_{q})^{p}Z_{q}\int_{x=2}^{\infty}\frac{1}{\log(x)^{2(p-1)}x}\mathrm{d}x=(Z_{\pi}/Z_{q})^{p}Z_{q}\int_{\log 2}^{\infty}\frac{1}{t^{2(p-1)}}\mathrm{d}t<\infty,$
as $p\geq 2$.
Now we estimate $\mathbb{P}_{X\sim\pi}(X>s)$ and $\mathbb{P}_{X\sim q}(X>s)$ for any $s>2$ respectively. For the former:
$\mathbb{P}_{X\sim\pi}\left(X>s\right)=Z_{\pi}\int_{s}^{\infty}\frac{1}{x^{p}}\mathrm{d}x=\frac{C_{1}}{s^{p-1}}.$
For the latter,
$\mathbb{P}_{X\sim q}\left(X>s\right)$ $=Z_{q}\int_{s}^{\infty}\frac{\log^{2}(x)}{x^{p+1}}\mathrm{d}x=Z_{q}\int_{1}^{\infty}\frac{\log^{2}(su)}{s^{p+1}u^{p+1}}\cdot s\mathrm{d}u$
$\leq\frac{Z_{q}}{s^{p}}\int_{1}^{\infty}\frac{2\log^{2}(s)+2\log^{2}(u)}{u^{p+1}}\mathrm{d}u$
$\leq\frac{C_{2}\log^{2}(s)}{s^{p}}+\frac{C_{3}}{s^{p}}\leq\frac{C_{4}\log^{2}(s)}{s^{p}},$
where the first inequality follows from $(a+b)^{2}\leq 2a^{2}+2b^{2}$.
Consider an IMH chain $(X_{t})_{t\geq 0}$ targeting $\pi$ with proposal $q$ starting at $x_{0}=3$. Fix any $t\geq 100$, define $A_{t}:=(t(\log t)^{3},\infty)$. Then the probability of $A_{t}$ under $\pi$ is
$\mathbb{P}_{X\sim\pi}\left(X\in A_{t}\right)=\frac{C_{1}}{t^{p-1}(\log t)^{3(p-1)}}.$
On the other hand, $X_{t}$ is in $A_{t}$ implies at least one of the proposals made at times $1,2,\ldots,t$ falls into $A_{t}$ (note that $x_{0}\notin A_{t}$ since $100(\log(100))^{3}\approx 10^{4}$). By the union bound, we have
$\mathbb{P}\left(X_{t}\in A_{t}\right)\leq t\cdot\mathbb{P}_{Y\sim q}\left(Y\in A_{t}\right)\leq t\cdot\frac{C_{4}\log^{2}(t(\log t)^{3})}{t^{p}(\log t)^{3p}}\leq\frac{C_{4}\log^{2}(t^{2})}{t^{p-1}(\log t)^{3p}}=\frac{4C_{4}(\log t)^{2}}{t^{p-1}(\log t)^{3p}},$
where the last inequality uses $\log(t)^{3}\leq t$ when $t\geq 100$. Therefore, we have the following lower bound on the TV distance
$\left|P^{t}(x_{0},\cdot)-\pi\right|_{\text{TV}}$ $\geq\mathbb{P}_{X\sim\pi}\left(X\in A_{t}\right)-\mathbb{P}\left(X_{t}\in A_{t}\right)$
$\geq\frac{C_{1}(\log t)^{3}}{t^{p-1}(\log t)^{3p}}-\frac{4C_{4}(\log t)^{2}}{t^{p-1}(\log t)^{3p}}.$
Since $(\log t)^{2}=o((\log t)^{3})$ as $t\rightarrow\infty$, there exists $t_{0}=t_{0}(p)$ and $C_{5}>0$ such that for any $t>t_{0}$:
$\left|P^{t}(x_{0},\cdot)-\pi\right|_{\text{TV}}$ $\geq\frac{C_{5}}{t^{p-1}(\log t)^{3(p-1)}}=\tilde{\Omega}(t^{-(p-1)}).$
A.5 Proofs of Section 5
A.5.1 Proof of Proposition 5.1
Proof.
Note that $\hat{F}_{u}$ is not bounded even if $|f|_{\infty}\leq 1$, because the sum in (36) can be arbitrarily large. By Minkowski’s inequality, for any $s\geq 1$,
$\mathbb{E}\left[|\hat{F}_{u}|^{s}\right]^{1/s}\leq\mathbb{E}\left[|\hat{F}(\mathbf{x}_{0})|^{s}\right]^{1/s}+\mathbb{E}\left[\left|\sum_{t=1}^{\tau-1}\{\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})\}\right|^{s}\right]^{1/s}.$ (101)
Furthermore, if $|f|_{\infty}\leq 1$ then $|\hat{F}(\mathbf{x})|\leq 1$ for all $\mathbf{x}$, thus
$\mathbb{E}\left[|\hat{F}_{u}|^{s}\right]^{1/s}\leq\mathbb{E}\left[|\hat{F}(\mathbf{x}_{0})|^{s}\right]^{1/s}+2\mathbb{E}\left[\mathds{1}(\tau>1)\left|\tau-1\right|^{s}\right]^{1/s}.$ (102)
Since $\hat{F}(\mathbf{x}_{0})\leq 1$ almost surely, $\mathbb{E}[|\hat{F}(\mathbf{x}_{0})|^{s}]^{1/s}$ is finite for all $s\geq 1$. The latter expectation is smaller than $\mathbb{E}[|\tau|^{s}]^{1/s}$. Thus, $\hat{F}_{u}$ has $s$ finite moments if $\tau$ has $s$ finite moments. Note that $\hat{F}_{u}$ can have higher moments as well: for example if $f$ is constant then $\hat{F}_{u}$ is constant.
Next, in order for $\tau$ to have $s\geq 1$ moments, we can resort to Proposition 4.3. If Assumption 2 holds with $p>s$, then $\mathbb{P}(\tau>t)\leq CN^{-1/2}t^{-p}$. We can then follow the proof of Proposition 8 in Douc et al. (2024), using Tonelli’s theorem:
$\mathbb{E}\left[\tau^{s}\right]$ $=\mathbb{E}\left[\int_{0}^{\infty}\mathds{1}(u<\tau)su^{s-1}\mathrm{d}u\right]$
$=\int_{0}^{\infty}su^{s-1}\mathbb{P}(\tau>u)\mathrm{d}u$
$=\sum_{i=0}^{\infty}\mathbb{P}(\tau>i)\int_{i}^{i+1}su^{s-1}\mathrm{d}u$
$\leq\sum_{i=0}^{\infty}\mathbb{P}(\tau>i)s(i+1)^{s-1}.$
The sum is finite under the assumption $p>s$. ∎
A.5.2 Proof of Proposition 5.2
Proof of Proposition 5.2.
We now consider the PIMH chain $(\mathbf{x}_{t})_{t\geq 0}$, started from $\bar{q}$. The case $t=0$ corresponds to Theorem 2.3. Let $t\geq 1$. We can assume that $f$ is non-negative, using the same separate treatment of $f_{+}$ and $f_{-}$ as in the beginning of the proof of Theorem 2.3.
We write
$\hat{F}^{\circ}:\mathbf{x}\mapsto\hat{F}(\mathbf{x})-\pi(f)=\frac{\sum_{n=1}^{N}\omega(x_{n})\{f(x_{n})-\pi(f)\}}{\sum_{m=1}^{N}\omega(x_{m})}.$ (103)
We can write
$\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[|\hat{F}(\mathbf{x}_{t})-\pi(f)|^{s}\right]$ $=\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}$
$=\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}\{\mathds{1}(A_{t})+\mathds{1}(A_{t}^{c})\},$
where the event $A_{t}$ represents “there was an acceptance in the first $t$ steps”.
In the event $A_{t}^{c}$, $\mathbf{x}_{t}=\mathbf{x}_{0}$ so
$\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}\cdot\mathds{1}(A_{t}^{c})$
$=\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{0})|^{s}\cdot\mathds{1}(A_{t}^{c})$
$\leq\int\bar{q}(\mathrm{d}\mathbf{x}_{0})|\hat{F}^{\circ}(\mathbf{x}_{0})|^{s},$
by bounding the indicator by one, and we can use Theorem 2.3 to obtain a bound in $N^{-s/2}$.
Now we consider the case $A_{t}$. For $1\leq j\leq t$ define the events
$A_{j,t}:=\{\mathbf{x}_{j-1}\neq\mathbf{x}_{j}=\mathbf{x}_{j+1}=\cdots=\mathbf{x}_{t}\},$
where $A_{j,t}$ is the event that there is a jump at time $j$ and no jump after that. Then $A_{j,t}\cap A_{j^{\prime},t}=\emptyset$ for $j\neq j^{\prime}$ and $A_{t}=\cup_{j=1}^{t}A_{j,t}$. We can decompose $\mathds{1}(A_{t})$ into $\sum_{j=1}^{t}\mathds{1}(A_{j,t})$ to get
$\int\bar{q}(\mathrm{d}\mathbf{x}_{0})P(\mathbf{x}_{0},\mathrm{d}\mathbf{x}_{1})\ldots P(\mathbf{x}_{t-1},\mathrm{d}\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}\mathds{1}(A_{t})$
$=\sum_{j=1}^{t}\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[|\hat{F}^{\circ}(\mathbf{x}_{t})|^{s}\mathds{1}(A_{j,t})\right]=\sum_{j=1}^{t}\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[\mathbb{E}\left\{\left.|\hat{F}^{\circ}(\mathbf{x}_{j})|^{s}\mathds{1}(A_{j,t})\right|\mathbf{x}_{j-1}\right\}\right].$
Conditional on $\mathbf{x}_{j-1}$,
$\int P(\mathbf{x}_{j-1},d\mathbf{x}_{j})P(\mathbf{x}_{j},d\mathbf{x}_{j+1})\cdots P(\mathbf{x}_{t-1},d\mathbf{x}_{t})|\hat{F}^{\circ}(\mathbf{x}_{j})|^{s}\mathds{1}\{\mathbf{x}_{j-1}\neq\mathbf{x}_{j}=\cdots=\mathbf{x}_{t}\}$
$=\int P(\mathbf{x}_{j-1},d\mathbf{x}_{j})|\hat{F}^{\circ}(\mathbf{x}_{j})|^{s}\mathds{1}\{\mathbf{x}_{j-1}\neq\mathbf{x}_{j}\}\int P(\mathbf{x}_{j},d\mathbf{x}_{j+1})\cdots P(\mathbf{x}_{t-1},d\mathbf{x}_{t})\mathds{1}\{\mathbf{x}_{j}=\cdots=\mathbf{x}_{t}\}$
$=\int P(\mathbf{x}_{j-1},d\mathbf{x}_{j})|\hat{F}^{\circ}(\mathbf{x}_{j})|^{s}\mathds{1}\{\mathbf{x}_{j-1}\neq\mathbf{x}_{j}\}r(\mathbf{x}_{j})^{t-j}$
$=\int\bar{q}(d\zeta)\alpha(\mathbf{x}_{j-1},\zeta)|\hat{F}^{\circ}(\zeta)|^{s}r(\zeta)^{t-j}.$
We can then upper bound $\alpha$ by one, and upper bound $\sum_{j=1}^{t}r(\zeta)^{t-j}$ by $(1-r(\zeta))^{-1}$ to obtain
$\sum_{j=1}^{t}\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q},\zeta\sim\bar{q}}\left[\alpha(\mathbf{x}_{j-1},\zeta)|\hat{F}^{\circ}(\zeta)|^{s}r(\zeta)^{t-j}\right]$
$\leq\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\sum_{j=1}^{t}r(\zeta)^{t-j}\right]$
$\leq\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\right].$
Next, split the expectation into the cases $\hat{Z}(\zeta)>2$ and $\hat{Z}(\zeta)\leq 2$:
$\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\right]$ $=\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)\leq 2)\right]+\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)>2)\right].$
When $\hat{Z}(\zeta)\leq 2$, since $r$ is increasing with $\hat{Z}$, we have $r(\zeta)\leq r(2)$ and thus $(1-r(\zeta))^{-1}\leq(1-r(2))^{-1}$. This yields:
$\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)\leq 2)\right]$ $\leq\frac{1}{1-r(2)}\int\bar{q}(\mathrm{d}\zeta)|\hat{F}^{\circ}(\zeta)|^{s}\mathds{1}(\hat{Z}(\zeta)\leq 2)$
$\leq\frac{1}{1-r(2)}\mathbb{E}_{\mathbf{x}_{0}\sim\bar{q}}\left[|\hat{F}^{\circ}(\mathbf{x}_{0})|^{s}\right].$
We obtain a bound in $N^{-s/2}$ using Theorem 2.3.
When $\hat{Z}(\zeta)>2$, from Lemma A.4, we have $r(\zeta)\leq 1-c_{p}(1/2)/(2\hat{Z}(\zeta))$. Thus $1-r(\zeta)\geq c_{p}(1/2)/(2\hat{Z}(\zeta))$, and $(1-r(\zeta))^{-1}\leq(2/c_{p}(1/2))\hat{Z}(\zeta)$. This yields:
$\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)>2)\right]\leq\frac{2}{c_{p}(1/2)}\int\bar{q}(\mathrm{d}\zeta)|\hat{F}^{\circ}(\zeta)|^{s}\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2).$
Since we assume $f\geq 0$, we can use the inequality (62):
$|\hat{F}^{\circ}(\zeta)|^{s}\leq\left(\max_{1\leq i\leq N}f(\zeta_{i})+q(\omega f)\right)^{s},$
from which we obtain
$\mathbb{E}_{\zeta\sim\bar{q}}$ $\left[|\hat{F}^{\circ}(\zeta)|^{s}\frac{1}{1-r(\zeta)}\mathds{1}(\hat{Z}(\zeta)>2)\right]$
$\leq\frac{2}{c_{p}(1/2)}\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})+q(\omega f)\right)^{s}\cdot\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]$
$\leq\frac{2^{s}}{c_{p}(1/2)}\left(\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{s}\cdot\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]+q(\omega f)^{s}\cdot\mathbb{E}_{\zeta\sim\bar{q}}\left[\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]\right).$
Using the facts that $\hat{Z}(\zeta)\leq 2(\hat{Z}(\zeta)-1)$ when $\hat{Z}(\zeta)>2$ and $\mathds{1}(\hat{Z}(\zeta)\geq 2)\leq|\hat{Z}(\zeta)-1|^{p-1}$, we obtain via Proposition 1.1:
$q(\omega f)^{s}\cdot\mathbb{E}_{\zeta\sim\bar{q}}\left[\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]$ $\leq 2q(\omega f)^{s}\mathbb{E}_{\zeta\sim\bar{q}}\left[|\hat{Z}(\zeta)-1|^{p}\right]$
$\leq\frac{M(p)}{N^{p/2}}q(\omega f)^{s}.$
For the remaining term, using Hölder’s inequality yields:
$\mathbb{E}_{\zeta\sim\bar{q}}$ $\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{s}\cdot\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]$
$\leq 2\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{r}\right]^{s/r}\cdot\mathbb{E}_{\zeta\sim\bar{q}}\left[\left|\hat{Z}(\zeta)-1\right|^{\frac{r}{r-s}}\mathds{1}(\hat{Z}(\zeta)>2)\right]^{1-s/r}.$
Under the assumptions, with $s\leq\frac{pr}{p+r+2}$, we have:
$\frac{r}{r-s}\leq\frac{p+r+2}{r+2}=1+\frac{p}{r+2}\leq p,$
where the inequality holds since $r\geq 2$ by assumption. This gives us:
$\mathbb{E}_{\zeta\sim\bar{q}}\left[\left|\hat{Z}(\zeta)-1\right|^{\frac{r}{r-s}}\mathds{1}(\hat{Z}(\zeta)>2)\right]^{1-s/r}$ $\leq\left(\frac{M(p)}{N^{p/2}}\right)^{1-s/r}.$
Finally we use the fact that $\max\{a_{1},\ldots,a_{n}\}\leq a_{1}+\cdots+a_{n}$ for non-negative $a_{i}$ to derive
$\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{r}\right]$
$=\mathbb{E}_{\zeta\sim\bar{q}}\left[\max_{1\leq i\leq N}f(\zeta_{i})^{r}\right]\leq\mathbb{E}_{x\sim q}\left[f(x)^{r}N\right],$
so that
$\mathbb{E}_{\zeta\sim\bar{q}}\left[\left(\max_{1\leq i\leq N}f(\zeta_{i})\right)^{s}\cdot\hat{Z}(\zeta)\mathds{1}(\hat{Z}(\zeta)>2)\right]$
$\leq 2\left(q(f^{r})N\right)^{s/r}\cdot\left(\frac{M(p)}{N^{p/2}}\right)^{1-s/r}.$
We end up with an exponent of $N$ equal to $s/r-(p/2)(r-s)/r$, which under the assumptions is less than $-s/2$, as detailed in the proof of Theorem 2.3. Therefore, we obtain an upper bound in $N^{-s/2}$ on all terms.
Remark A.1.
Under Assumption 1, PIMH converges in total variation. Thus, $(\mathbf{x}_{t})$ converges weakly to $\bar{\pi}$. We consider the transformation $\mathbf{x}\mapsto|\hat{F}^{\circ}(\mathbf{x})|^{q}$ and Fatou’s lemma as in Theorem 3.4 of Billingsley (1999), to obtain
$\mathbb{E}_{\bar{\pi}}[|\hat{F}^{\circ}(\mathbf{x})|^{q}]\leq\lim\inf_{t}\mathbb{E}[|\hat{F}^{\circ}(\mathbf{x}_{t})|^{q}].$
Thus, the bound of Proposition 5.2, valid for all $t\geq 0$, applies also to the $s$-th moment of $\hat{F}(\mathbf{x})-\pi(f)$ under $\bar{\pi}$.
A.5.3 Proof of Proposition 5.3
Proof of Proposition 5.3.
We start as in the proof of Proposition 5.1 in Appendix A.5.1, and employ Theorem 2.3 for the moments of the error of IS with unbounded functions. Regarding the bias cancellation term,
BC $=\sum_{t=1}^{\infty}\Delta_{t}\mathds{1}(\tau>t),$ (104)
we use Minkowski with exponent $s\geq 1$:
$\mathbb{E}\left[|\text{BC}|^{s}\right]^{1/s}$ $\leq\sum_{t=1}^{\infty}\mathbb{E}\left[|\Delta_{t}|^{s}\mathds{1}(\tau>t)\right]^{1/s}.$ (105)
Next, for each time $t$, using Hölder’s inequality with an arbitrary $\kappa>1$,
$\mathbb{E}\left[|\Delta_{t}|^{s}\mathds{1}(\tau>t)\right]\leq\mathbb{E}\left[|\Delta_{t}|^{s\kappa}\right]^{1/\kappa}\mathbb{P}(\tau>t)^{(\kappa-1)/\kappa}.$ (106)
For the sum over $t$ in (105) to be finite, and using Proposition 4.3 to bound $\mathbb{P}(\tau>t)$, we have the condition on $\kappa$ and $s$,
$-\frac{p(\kappa-1)}{s\kappa}<-1\quad\Leftrightarrow\quad\kappa>p/(p-s).$
To establish the finiteness of $\mathbb{E}\left[|\Delta_{t}|^{s\kappa}\right]$ we can resort to Proposition 5.2 if $s\kappa$ satisfies the condition
$s\kappa\leq\frac{pr}{p+r+2}.$
We can find such $\kappa$ if
$\frac{ps}{p-s}<\frac{pr}{p+r+2}.$
A.5.4 Proof of Proposition 5.4
Proof of Proposition 5.4.
We follow the proof of Proposition 5.3, with $s=2$. We thus have a exponent $\kappa>1$ that must satisfy $\kappa>p/(p-2)$, and $2\kappa\leq pr/(p+r+2)$. We choose any number $\kappa$ strictly between $p/(p-2)$ and $pr/(2p+2r+4)$, which is possible by assumption, since
$1<\frac{p}{p-2}<\frac{pr}{2p+2r+4}\Leftrightarrow 2p+4r+4<rp.$
For that $\kappa$, we can apply Proposition 5.2 to bound $\mathbb{E}[|\Delta_{t}|^{2\kappa}]^{1/\kappa}$ by a constant times $N^{-1}$. Meanwhile, the sum $\sum_{t=1}^{\infty}\mathbb{P}(\tau>t)^{(\kappa-1)/(\kappa s)}$ is finite using Proposition 4.3, and is of the form $CN^{-a}$ for some positive $a$, namely $a=(\kappa-1)/(2\kappa s)$. Thus $\mathbb{E}{|\text{BC}|^{2}}$ can be bounded by a constant times $N^{-1-a}$ for some positive $a$, and finds itself negligible in front of the MSE of IS as $N\to\infty$. ∎